### v6

# Huber-robust likelihood ratio tests for composite nulls and alternatives

Abstract
We propose an e-value based framework for testing arbitrary composite nulls against composite alternatives, when an $\epsilon$ fraction of the data can be arbitrarily corrupted. Our tests are inherently sequential, being valid at arbitrary data-dependent stopping times, but they are new even for fixed sample sizes, giving type-I error control without any regularity conditions. We first prove that least favourable distribution (LFD) pairs yield optimal e-values for testing arbitrary composite nulls against composite alternatives. Then we show that if an LFD pair exists for some composite null and alternative, then the LFDs of Huber’s $\epsilon$-contamination neighborhoods around that pair form the optimal LFD pair for the corresponding robustified hypotheses. Furthermore, where LFDs do not exist, we develop new robust composite tests for general settings. Our test statistics are a nonnegative supermartingale under the (robust) null, even under a sequentially adaptive (non-i.i.d.) contamination model where the conditional distribution of each observation given the past data lies within an $\epsilon$ total variation (TV) ball of some distribution in the original composite null. When LFDs exist, our supermartingale grows to infinity exponentially fast under any distribution in the ($\epsilon$ TV-corruption of the) alternative at the optimal rate. When LFDs do not exist, we provide an asymptotic growth rate analysis, showing that as $\epsilon\to 0$, the exponent converges to the corresponding Kullback–Leibler divergence, recovering the classical optimal non-robust rate. Simulations validate the theory and demonstrate reasonable practical performance.
1 Introduction
In statistical hypothesis testing, the assumption that hypothesized models are perfectly specified is often far from reality. Real-world data rarely conforms perfectly to our idealized models, making it crucial to develop robust testing methodologies that can withstand small — and potentially adversarial — deviations from the idealized models.
Let $\mathcal{M}$ denote the set of all probability distributions over some measurable space $(\Omega,\mathcal{A})$. We consider both batch and sequential tests, but using the latter as a way of getting to the former. In the former setting, we observe a batch of data $X_{1},\dots,X_{n}$ from some unknown distribution $Q\in\mathcal{M}$. In the latter setting, we sequentially observe data points $X_{1},X_{2},\cdots$ from $Q$.
Given $\alpha\in(0,1)$, this paper will consider the general problem of designing a level-$\alpha$ test for the null hypothesis $Q\in H_{0}$ vs. the alternative hypothesis that $Q\in H_{1}$, for some given sets $H_{0},H_{1}\subset\mathcal{M}$, when an $\epsilon$ fraction of the data can be arbitrarily corrupted by an adversary or, more generally, when the true data distribution lies within an $\epsilon$ neighborhood of the hypothesized models. Our tests are valid (control type-1 error at $\alpha$) at arbitrary stopping times; these yield batch tests at fixed times as a special case.
Our advances build on two threads of the literature: classic papers in robustness and recent advances in composite null testing using e-values. First, for a simple point null and alternative, Huber [1] constructed a robust sequential likelihood ratio test. This will be a central starting point for the current paper. While Huber constructed least favourable distribution (LFD) pair to design optimal tests in both finite sample size and sequential settings, he did not provide a way to construct “power-one” sequential tests that can control the type-1 error at a prespecified level $\alpha$.
The second thread of literature that is relevant involves recent fundamental advances in (non-robust) composite hypothesis testing. For example, Grünwald et al. [2] introduces growth-rate optimal (GRO) e-variables for testing problems with composite null and alternative. The universal inference method [3] used an e-value to propose a level-$\alpha$ test for any composite null without requiring any regularity conditions, but requiring reference measures. Recently, Larsson et al. [4] constructed the log-optimal numeraire e-value based on the Reverse Information Projection (RIPr), which also does not require any regularity conditions or even reference measures, and is always more powerful than universal inference. We will introduce the history and definition of the RIPr in detail later, but it has been of great interest in information theory, and more recently, statistics, for several decades [5, 2, 6]. Since the numeraire always dominates the universal inference e-value in the non-robust setting, we only extend the numeraire to the robust setting. We use the robust numeraire to construct a nonnegative supermartingale under the (contaminated) null, that is easy to threshold to get a level-$\alpha$ test. While the finite-sample optimality of our methods is unresolved, our supermartingales are, at least in a certain weak limiting sense, proven to be log-optimal, a notion of optimality that has been long employed in information theory [7, 8].
To summarize, this work combines and significantly extends the techniques in [1, 2] and [4] to yield new fixed-sample and sequential robust tests for general composite nulls and alternatives. Beyond testing, the proposed supermartingales can also directly serve as building blocks for Huber-robust change detection [9] and localization [10].
The rest of this introduction recaps Huber’s idea, introduces e-values and test supermartingales, discusses related work in more detail and delineates our contributions relative to these.
1.1 Huber’s proposal for a simple null versus simple alternative and our modification
Consider sequentially testing $Q=P_{0}$ versus $Q=P_{1}$, for some given $P_{0},P_{1}\in\mathcal{M}$. By this, we mean that we observe a stream of data from $P_{0}^{\infty}$ or from $P_{1}^{\infty}$. Let $p_{0}$ and $p_{1}$ be the respective densities with respect to some dominating measure $\mu$. The classical sequential probability ratio test (by Wald) for this testing problem is not robust: a single factor $p_{1}(X_{j})/p_{0}(X_{j})$ equal or close to 0 or $\infty$ may ruin the entire nonnegative martingale $T_{n}=\prod_{i=1}^{n}\frac{p_{1}(X_{i})}{p_{0}(X_{i})}$.
Huber [1] formalized the problem of robustly testing a simple $P_{0}$ against a simple $P_{1}$ by assuming that the true underlying distribution $Q$ lies in some appropriate neighborhood of either of the idealized models $P_{0}$ or $P_{1}$. To account for the possibility of small deviations from the idealized models $P_{j}$, he expanded the hypotheses to:
$H^{\epsilon}_{j}=\{Q\in\mathcal{M}:Q=(1-\epsilon)P_{j}+\epsilon H,H\in\mathcal{M}\},$ (1)
or, more generally, to
$H^{\epsilon}_{j}=\{Q\in\mathcal{M}:D_{\operatorname{TV}}(P_{j},Q)\leqslant\epsilon\},$ (2)
where $j=0,1$ and $D_{\operatorname{TV}}$ denotes the total variation distance. Thus, a robust test of $P_{0}$ versus $P_{1}$ is effectively a test of the composite null $H^{\epsilon}_{0}$ against the composite alternative $H^{\epsilon}_{1}$. Note that (2) is strictly larger than (1). He proposed risk-optimal tests $\phi$ for the following three notions of optimality:
$\min_{\phi}\max_{j=0,1}\sup_{Q_{j}\in H^{\epsilon}_{j}}R(Q_{i},\phi)$ (3)
$\min_{\phi}\sup_{Q_{1}\in H^{\epsilon}_{1}}R(Q_{1},\phi)\;\;\text{s.t. }\sup_{Q_{0}\in H^{\epsilon}_{0}}R(v_{0},\phi)\leqslant\alpha$
$\min_{\phi}\sup_{(Q_{0},Q_{1})\in H^{\epsilon}_{0}\times H^{\epsilon}_{1}}\big[\pi_{0}R(Q_{0},\phi)+\pi_{1}R(Q_{1},\phi)\big].$
He solved the preceding problems by looking for a “least favorable distribution pair” in terms of risk, where the risk $R$ for a hypothesis test $\phi$ is defined as
$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }H_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }H_{0}).$ (4)
The notion of a least favorable distribution pair in terms of risk is formally defined below.
Definition 1.1.
A pair $(P_{0}^{*},P_{1}^{*})\in\mathcal{P}_{0}\times\mathcal{P}_{1}$ is called least favorable distribution (LFD) pair in terms of risk $R$ for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, if for every likelihood ratio test $\phi$ between $P_{0}^{*}$ and $P_{1}^{*}$,
$R(P_{j}^{*},\phi)\geq R(P_{j},\phi),\penalty 10000\ \text{ for all }P_{j}\in\mathcal{P}_{j},\penalty 10000\ j=0,1.$
Huber derived the LFD pair $(Q_{0,\epsilon},Q_{1,\epsilon})$ in $H^{\epsilon}_{0}\times H^{\epsilon}_{1}$, such that any probability ratio test of $Q_{0,\epsilon}$ vs. $Q_{1,\epsilon}$ is precisely the optimal test for $H^{\epsilon}_{0}$ vs. $H^{\epsilon}_{1}$.
He showed that their likelihood ratio $\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}=\max\{c^{\prime},\min\{c^{\prime\prime},\frac{p_{1}(X)}{p_{0}(X)}\}\}$ is a truncation of the original likelihood ratio $\frac{p_{1}(X)}{p_{0}(X)}$, at some specific values of the $c^{\prime}$ and $c^{\prime\prime}$, which depend on $\epsilon,P_{0}$ and $P_{1}$. And hence, Huber’s truncated likelihood ratio test is the optimal solution for the finite-sample problems defined in (LABEL:eq:risk-optimality). He also considered (two-sided) sequential tests with the stopping time defined below:
$N=\inf\left\{n\geq 0:S_{n}\leqslant a\text{ or }S_{n}\geq b\right\},$ (5)
where $a<b$ are fixed numbers. The testing procedure is to stop at stage $N$ and reject $H^{\epsilon}_{0}$ if $S_{n}\geq b$ and accept $H^{\epsilon}_{0}$ if $S_{n}\leqslant a$. Huber [1] proved that $Q_{j,\epsilon}\in H_{j}^{\epsilon},j=0,1$ are “least favorable distribution pair”in the sequential setting as well, i.e,
$Q_{0,\epsilon}[S_{n}\geq b]=\sup\{Q[S_{n}\geq b]:Q\in H_{0}^{\epsilon}\},\text{ and }$ (6)
$Q_{1,\epsilon}[S_{n}\leqslant a]=\sup\{Q[S_{n}\leqslant a]:Q\in H_{1}^{\epsilon}\}.$ (7)
and hence, the test is optimal for type-I and type-II error probabilities.
Our level-$\alpha$, anytime-valid variant of Huber’s test: Huber does not describe how exactly to calculate the thresholds if a targeted type-I error of $\alpha$ is specified. We prove an interesting general result (that holds true even outside the context of this paper) that the likelihood ratio of the LFD pair is actually an e-value for the composite null $H_{0}^{\epsilon}$. Further, it can also be shown to be growth rate optimal (log-optimal). That allows us to construct a level-$\alpha$ test, which has power one. As a side benefit, it will also be an “anytime-valid” test [11] for this problem, to be defined later. We will actually construct an “anytime-valid e-value” (or an e-process, to be defined later) that is a continuous measure of evidence against the null, which can be thresholded at $1/\alpha$ to yield a level-$\alpha$ test. Though we do not expand on the following point further, we note in passing that this e-process can also be converted to an “anytime-valid p-value”, which would allow us to continuously monitor the data throughout the experiment, report a p-value at any stopping time, and also reject the null at a level $\alpha$ when that p-value drops below $\alpha$. A fixed sample size test can be obtained by simply stopping monitoring at a fixed time $n$.
This framework is an offshoot of Robbins’ power-one sequential tests (or one-sided sequential probability ratio tests). The framework is rooted in the construction of “test supermartingales” [12] — or, more generally, “e-processes” [13] — which can be interpreted as the wealth of a gambler playing a stochastic game; these are introduced below. To achieve anytime-valid inference, we construct a test supermartingale under $H^{\epsilon}_{0}$, ensuring its validity at arbitrary data-dependent stopping times, accommodating continuous monitoring and analysis of accumulating data, and optional stopping or continuation. In later sections, we extend our method to composite nulls and composite alternatives, leveraging the techniques of “predicable plug-in” and “Reverse Information Projection”, which are well-known yet sophisticated tools employed in the non-robust setting.
Before we delve into the details of our method, it is crucial to discuss what test (super)martingales are and how they play a key role in constructing sequential anytime-valid tests.
1.2 Background on e-values, test supermartingales, one-sided tests, growth rate and consistency
An e-variable is a nonnegative random variable that has expectation at most one under the null. Mathematically, a nonnegative random variable $B$ is an e-variable for the null $P\in H_{0}^{\epsilon}$ if $\mathbb{E}_{P}(B)\leqslant 1$, for all $P\in H_{0}^{\epsilon}$. The value realized by an e-variable (that is, its instantiation) is called an e-value.
An integrable process $M\equiv\{M_{n}\}$ that is adapted to a filtration $\mathscr{F}\equiv\{\mathscr{F}_{n}\}_{n\geq 0}$, is called a martingale for $H_{0}^{\epsilon}$ if
$\mathbb{E}_{{P}}[M_{n}\mid\mathscr{F}_{n-1}]=M_{n-1},\forall P\in H_{0}^{\epsilon}$ (8)
for all $n\geq 1$. $M$ is called a supermartingale for $H_{0}^{\epsilon}$ if for all $n\geq 1$,
$\mathbb{E}_{{P}}[M_{n}\mid\mathscr{F}_{n-1}]\leqslant M_{n-1},\forall P\in H_{0}^{\epsilon}.$ (9)
Crucially, $M$ is called a test (super)martingale for $H_{0}^{\epsilon}$ if it is a (super)martingale for every $P\in H_{0}^{\epsilon}$, and if it is nonnegative with $M_{0}=1$. A stopping time $\tau$ is a nonnegative integer-valued random variable such that $\{\tau\leqslant n\}\in\mathscr{F}_{n}$ for each $n\in\mathbb{N}$. Denote by $\mathcal{T}$ the set of all stopping times, including ones that may never stop.
Ville’s inequality [14] implies that the test (super)martingale satisfies
$\sup_{P\in H_{0}^{\epsilon}}P(\exists n\in\mathbb{N}:M_{n}\geq 1/\alpha)\leqslant\alpha.$ (10)
See [15, Lemma 1] for a short proof. The above equation is equivalent to $P(M_{\tau}\geq 1/\alpha)\leqslant\alpha,\forall\tau\in\mathcal{T},P\in H_{0}^{\epsilon}$; see [16, Lemma 3]. This ensures that if we stop and reject the null at the stopping time
$\tau_{\alpha}=\inf\left\{n\geq 1:M_{n}\geq 1/{\alpha}\right\},$ (11)
it results in a level-$\alpha$ sequential test, meaning that if the null is true, the probability that it ever stops falsely rejecting the null is at most $\alpha$ (under the null, $\tau_{\alpha}=\infty$ with probability $1-\alpha$).
The above definition of a sequential test differs from Wald’s original ideas [17]. In the latter, the null hypothesis might eventually be accepted or rejected, with a predetermined stopping rule based on the desired Type-I and Type-II error rates. In contrast, our framework aligns with Robbins’ “power-one tests” [18, 19], or one-sided tests, where one only specifies a target Type-I error level and we only stop for rejecting the null, but never stop for acceptance. Such a test is called consistent if it is (asymptotically) power one under any alternative $P\in H^{1}_{\epsilon}$.
A test (super)martingale is called consistent (or power one) for $H_{1}^{\epsilon}$, if $\lim_{n\to\infty}M_{n}=\infty$ almost surely for any $P\in H_{1}^{\epsilon}$, meaning that under the alternative, it accumulates infinite evidence against the null in the limit and eventually crosses any finite threshold for rejection. As Kelly noted in his seminal work [7], the evidence grows exponentially fast in i.i.d. settings, so one can define the “growth rate” of $M$ is defined as $\inf_{P\in H_{1}^{\epsilon}}\liminf_{n\to\infty}\log M_{n}/n$ [20, 2, 21]. A positive growth rate implies consistency. The test (super)martingale with the largest possible growth rate is called log-optimal, an optimality notion advocated by many influential researchers, like [22], [8], [20] and [2] amongst others.
If $\epsilon=0$ and $H_{0}=\{P_{0}\},H_{1}=\{P_{1}\}$, then we are back to the standard non-robust testing of $P_{0}$ against $P_{1}$ (except that we desire a one-sided, or power-one, sequential test). In this setting,
$T_{n}=\prod_{i=1}^{n}\frac{p_{1}(X_{i})}{p_{0}(X_{i})}$ (12)
is a test martingale for $P_{0}$. Recalling (11), we may reject the null at the stopping time $\inf\left\{n\geq 1:T_{n}\geq 1/{\alpha}\right\}$ to get a level $\alpha$ sequential test. The growth rate of the test reduces to $\mathbb{E}_{P_{1}}\log\frac{p_{1}(X)}{p_{0}(X)}=D_{\operatorname{KL}}(P_{1},P_{0})$. And it is the optimal growth rate for testing $P_{0}$ vs. $P_{1}$ [20], a result with roots in [7] and [22].
1.3 Related Work
The formalization of robust statistics can be traced back to the mid-20th century. Early pioneers in the development of the concept include [23, 1, 24]. Prominent ideas in robust estimation include, among others, M-estimators, trimming, and influence functions. The use of M-estimators in robust statistical procedures dates back to [25], which achieves robustness by curbing and bounding the influence that individual data points can make on the statistics. On the other hand, trimming refers to the practice of directly discarding outliers [26], and trimmed means have long been known to be robust [27]. More recent developments extend this framework by incorporating data-driven methods, using Wasserstein, MMD, and entropic-regularized Wasserstein distances. For example, [28] constructs uncertainty sets under the null and the alternative distributions, which are sets centered around the empirical distribution defined via the Wasserstein metric. [29] proposes a data-driven approach to robust hypothesis testing using Sinkhorn uncertainty sets. [30] uses moment-constrained uncertainty sets. [31] introduces a kernel-based approach to robust hypothesis testing, where they construct tests that optimize worst-case performance within an MMD uncertainty set. The main difference of our work from all of these is that we consider (a generalization of) Huber’s original $\epsilon$-contamination model as our notion of robustness, which is akin to using the TV distance.
Recently, [32] proposed a robust version of the universal inference [3] approach for constructing valid confidence sets under weak regularity conditions, despite possible model misspecification. However, their notion of robustness differs from ours, and they test whether $P_{0}$ or $P_{1}$ is closer to the true data-generating distribution, meaning that before contamination it is a simple versus simple hypothesis test while we consider composite classes (even before contamination). [33] introduced Huber-robust confidence sequences leveraging supermartingales. The work of [34] is related to ours, as they investigated the multi-armed bandit problem under the Huber corruption model and briefly discussed its connection to the mean testing problem in their Appendix. They analyzed the minimizer of a function involving KL divergences, which is related to RIPr used in our methodology. However, their analysis is restricted to the Gaussian setting with known variance, whereas our framework and theory are general. Furthermore, their work addresses only the $\epsilon$ corruption model (1), while our approach accommodates the more general settings including $\epsilon$-TV neighborhoods (2).
Sequential hypothesis testing has a long-standing history, beginning with the sequential probability ratio test (SPRT) of [17]. There have been a few studies in the literature that address the feasibility of robustifying sequential tests, mostly notably the works by [1] and by [35], both of them robustifying the likelihood ratio by censoring. [35] considered the sequential testing of two distributions $P_{-\epsilon}$ and $P_{\epsilon}$, where these two distributions approach each other as $\epsilon\to 0$ and proved (under regularity assumptions) that the SPRT based on the least favorable pair of distributions given by [1] is asymptotically least favorable for the expected sample size and is asymptotically minimax. While these earlier tests yield valid inference at a particular prespecified stopping rule, e-values and e-processes (as used in the current paper) have recently emerged as a promising tool to construct “anytime-valid” tests [11]. Further, these previous works only considered simple versus simple tests (when contamination is excluded), while we work on composite tests.
1.4 Contributions
We first show in general that the likelihood ratio of an LFD pair is an e-value, which is also growth rate optimal (or, log optimal). As a consequence of the result, it follows that Huber’s truncated likelihood ratio test statistic [1] is in fact already a test supermartingale for $H_{0}^{\epsilon}$. The supermartingale tools enable us to generalize the usual Huber contamination model, demonstrating that our sequential test retains type-I error validity even under sequentially adaptive contaminations under the null, as well as at arbitrary stopping times, thereby extending Huber’s contributions in this case.
Next, we extend the framework to handle both composite nulls and composite alternative hypotheses (before contamination), a direction that, to the best of our knowledge, has not been pursued previously. We prove that when an LFD pair exists for a composite null and alternative, the LFDs of Huber’s $\epsilon$-contamination neighborhoods around that pair form the optimal LFD pair for the corresponding robustified hypotheses. Together with the earlier result, this enables the construction of optimal tests across all regimes: fixed-sample settings, two-sided (Wald-type) sequential settings, and anytime-valid testing scenarios.
Furthermore, where an LFD pair does not exist for the pre-contamination composite null and composite alternative, we develop new robust composite tests for general settings, where we utilize very recent advances in composite null hypothesis testing, such as universal inference [3] and the reverse information projection [2, 6, 4]. However, for this case, while we cannot prove finite sample log-optimality, we do analyze the growth rate under the alternative, which, as $\epsilon\to 0$, converges to the optimal non-robust growth rate (a certain KL divergence between null and alternative), under some assumptions. To justify these assumptions, we provide sufficient conditions under which they hold. We emphasize that our robust tests control the type-I error at level $\alpha$ without requiring any assumptions on the pre-contamination composite null and composite alternative.
To summarize, this paper provides a relatively comprehensive set of methods for Huber-robust sequential hypothesis testing, with a theory that accurately predicts performance in experiments.
1.5 Paper outline
The rest of the paper is organized as follows. In Section 2, we show that the likelihood ratio of the LFD pair is an e-value, and in fact the log-optimal e-value. In Section 3, we prove that if an LFD pair exists for some composite null and alternative, then the LFDs of Huber’s $\epsilon$-contamination neighborhoods around that pair form the optimal LFD pair for the corresponding robustified hypotheses, yielding log-optimal e-values for this case as well. Finally, in Section 4, we consider general composite nulls and alternatives for which an LFD pair may not exist, and fixed $\epsilon$ optimality for tractable methods appears hard to prove, and demonstrate that our methods are at least asymptotically optimal as $\epsilon\to 0$. Section 5 presents a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical performance of our approach. This article is concluded in Section 6. Omitted proofs of the theoretical results are provided in Appendix A.
2 LFD pairs yield log-optimal e-values
In this section, we prove a very general result that if a least favorable distribution pair for testing any two hypotheses $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ exists, then their likelihood ratio yields optimal e-values. The implications are that the least favorable distribution pair that [1] defined for solving (LABEL:eq:risk-optimality) also solves the problem of finding “growth rate optimal in worst-case (GROW)” or “log optimal” e-values. [2] introduced the latter notion and showed their existence under several assumptions, including full support of distributions and convexity of $\mathcal{P}_{0},\mathcal{P}_{1}$. Our theorem does not require such assumptions (but needs an LFD pair).
Theorem 2.1.
If $(P_{0}^{*},P_{1}^{*})$ is a least favorable distribution (LFD) pair in terms of the risk defined in (4) for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, then $\frac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}$ is a GROW e-variable for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$. That is, if $\mathcal{E}(\mathcal{P}_{0})$ denotes the set of all e-variables for $\mathcal{P}_{0}$, we have that $\frac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\in\mathcal{E}(\mathcal{P}_{0})$ and
$\sup_{E\in\mathcal{E}(\mathcal{P}_{0})}\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log E\right)=\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})=\inf_{(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}}D_{\operatorname{KL}}(P_{1},P_{0}).$
Proof.
Since $(P_{0}^{*},P_{1}^{*})$ is an LFD pair for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, for any $P_{0}\in\mathcal{P}_{0},$ we have
${R(P_{0},\phi)\leqslant R(P_{0}^{*},\phi)}$
[…]
${\implies{P_{0}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)>\eta\right)\leqslant{P_{0}^{*}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)>\eta\right),\text{ for all }\eta}$
${\implies\int_{\eta\geq 0}{P_{0}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)>\eta\right)\,d\eta\leqslant\int_{\eta\geq 0}{P_{0}^{*}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)>\eta\right)\,d\eta}$
${\implies\mathbb{E}_{P_{0}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)\right)\leqslant\mathbb{E}_{P_{0}^{*}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)\right)=1.}$
Therefore, $\frac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}$ is an e-variable. Now, for the log-optimality part, we first show that $\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)\geq\mathbb{E}_{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)$. Since $(P_{0}^{*},P_{1}^{*})$ is the least favorable pair, for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, for any $P_{1}\in\mathcal{P}_{1}$,
${R(P_{1},\phi)\leqslant R(P_{1}^{*},\phi)}$
[…]
${\implies{P_{1}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)\leqslant\eta\right)\leqslant{P_{1}^{*}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)\leqslant\eta\right),\text{ for all }\eta}$
${\implies{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>\log\eta\right)\,\geq{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>\log\eta\right)\ }\text{ for all }\eta>0$
${\implies\int_{t\geq 0}{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt\geq\int_{t\geq 0}{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt}.$
Similarly, for any $P_{1}\in\mathcal{P}_{1}$,
${R(P_{1},\phi)\leqslant R(P_{1}^{*},\phi)}$
[…]
${\implies{P_{1}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)<\eta\right)\leqslant{P_{1}^{*}}\!\left(\tfrac{dP_{1}^{*}}{dP_{0}^{*}}(X)<\eta\right),\text{ for all }\eta}$
${\implies{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}<\log\eta\right)\,\leqslant{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}<\log\eta\right)\ }\text{ for all }\eta>0$
${\implies\int_{t\geq 0}{P_{1}}\!\left(-\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt\leqslant\int_{t\geq 0}{P_{1}^{*}}\!\left(-\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt}.$
Therefore, for any $P_{1}\in\mathcal{P}_{1}$,
$\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)$ $=\int_{t\geq 0}{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt-\int_{t\geq 0}{P_{1}}\!\left(-\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt$
$\geq\int_{t\geq 0}{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt-\int_{t\geq 0}{P_{1}^{*}}\!\left(-\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}>t\right)\,dt$
$=\mathbb{E}_{P_{1}^{*}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*}).$
So, we have shown above that $\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*}).$ Now, suppose, if possible, $E$ is an e-varible for $\mathcal{P}_{0}$, which satisfies $\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log E\right)>D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*}).$ Then, we would have $\mathbb{E}_{P_{1}^{*}}\!\left(\log E\right)>D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})$, this is a contradiction because $E$ is an e-variable for the simple null ${P}_{0}^{*}$ (against simple alternative $P_{1}^{*}$) as well and hence its maximum possible growth rate is $D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})$. Thus, we have proved $\sup_{E\in\mathcal{E}(\mathcal{P}_{0})}\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log E\right)=\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*}).$
If $D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})>\inf_{(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}}D_{\operatorname{KL}}(P_{1},P_{0})$, then $D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})>D_{\operatorname{KL}}(P_{1},P_{0})$, for some $(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}$, but then $dP_{1}^{*}/dP_{0}^{*}$ is an e-value for $P_{0}$ vs $P_{1}$, whose growth rate $\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)\geq\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})>D_{\operatorname{KL}}(P_{1},P_{0})$, but $D_{\operatorname{KL}}(P_{1},P_{0})$ is largest possible growth rate for an e-value for $P_{0}$ vs $P_{1}$, which is a contradiction.
Therefore, we finally have
$\sup_{E\in\mathcal{E}(\mathcal{P}_{0})}\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log E\right)=\inf_{P_{1}\in\mathcal{P}_{1}}\mathbb{E}_{P_{1}}\!\left(\log\tfrac{dP_{1}^{*}(X)}{dP_{0}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})=\inf_{(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}}D_{\operatorname{KL}}(P_{1},P_{0}).$
Remark 2.2.
The last part of the theorem $D_{\operatorname{KL}}(P_{1}^{*},P_{0}^{*})=\inf_{(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}}D_{\operatorname{KL}}(P_{1},P_{0})$ also follows from Corollary 1 of [36], but we have a direct proof here.
Remark 2.3.
It is easy to verify that LFD is symmetric around the null and alternative hypotheses. That is, $(P_{0}^{*},P_{1}^{*})$ is an LFD pair in terms of the risk defined in (4) for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ is equivalent to saying $(P_{1}^{*},P_{0}^{*})$ is an LFD pair in terms of the risk defined in (4) for testing $\mathcal{P}_{1}$ vs. $\mathcal{P}_{0}$. And hence, we also have that $\frac{dP_{0}^{*}(X)}{dP_{1}^{*}(X)}$ is a GROW e-variable for testing $\mathcal{P}_{1}$ vs. $\mathcal{P}_{0}$, that is, $\frac{dP_{0}^{*}(X)}{dP_{1}^{*}(X)}\in\mathcal{E}(\mathcal{P}_{1})$ and
$\sup_{E\in\mathcal{E}(\mathcal{P}_{1})}\inf_{P_{0}\in\mathcal{P}_{0}}\mathbb{E}_{P_{0}}\!\left(\log E\right)=\inf_{P_{0}\in\mathcal{P}_{0}}\mathbb{E}_{P_{0}}\!\left(\log\tfrac{dP_{0}^{*}(X)}{dP_{1}^{*}(X)}\right)=D_{\operatorname{KL}}(P_{0}^{*},P_{1}^{*})=\inf_{(P_{0},P_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}}D_{\operatorname{KL}}(P_{0},P_{1}),$
where $\mathcal{E}(\mathcal{P}_{1})$ is the set of all e-variables for $\mathcal{P}_{1}.$
The above theorem allows us to construct level-$\alpha$ tests using Huber’s LFD pairs for testing (1) or (2), which we elaborate on next. However, it is worth noting that the applicability of this theorem extends beyond these specific settings, for instance, to classical testing problems such as those involving monotone likelihood ratio (MLR) families or the $2\times 2$ contingency table. Interestingly, [2] identified GROW e-variables in these latter settings, although their connection to LFDs was not explicitly recognized.
2.1 An adaptive contamination model for the null $H^{\epsilon,\infty}_{0}$ and alternative $H^{\epsilon,\infty}_{1}$
Our test retains its validity properties under the null hypothesis and growth rate under the alternative, even in the presence of an adaptive contamination model, where the data sequence $X_{1},X_{2},\cdots$ is generated such that the conditional distribution of $X_{n}$ given $X^{n-1}$ lies within $H^{\epsilon}_{i}$, where $H^{\epsilon}_{i}$ is some $\epsilon$ neighborhood around the hypothesized distribution $P_{i}$, for example, $\epsilon$-contamination or total variation balls defined in (1) or (2). We denote the set of all possible joint distributions of the sequence $X_{1},X_{2},\cdots$ under the adaptive contamination model as $H^{\epsilon,\infty}_{i}$. In the special case where the distribution of each $X_{n}$ is independent of the previously observed data, then
$H^{\epsilon,\infty}_{i}=H^{\epsilon}_{i}\times H^{\epsilon}_{i}\times\cdots,$ (13)
but we emphasize that $H^{\epsilon,\infty}_{i}$ allows for non-i.i.d. sequentially adaptive adversarial contamination. In words, while $P_{i}$ is fixed and known, the unknown contaminations may vary over time, and indeed, the contamination at each time $n$ can be influenced by the previously observed data. Formally, at each $n$, the adversary defines a measurable mapping $Q_{n}:\Omega^{n-1}\to H^{\epsilon}_{i}$ and $X_{n}$ is drawn from the distribution $Q_{n}(x^{n-1}),$ given that $X^{n-1}=x^{n-1}$; here we use the notation $a^{n-1}=(a_{1},\dots,a_{n-1})$ and $x^{n-1}$ denotes the realized value of the random vector $X^{n-1}$.
2.2 Robust test for simple null vs. simple alternative
[1] defined the LFD pair $(Q_{0,\epsilon},Q_{1,\epsilon})$ for (1) by their densities $(q_{0,\epsilon},q_{1,\epsilon})$ as follows:
[…]
The numbers $0\leqslant c^{\prime}<c^{\prime\prime}\leqslant\infty$ are constants that are determined to ensure $q_{0,\epsilon},q_{1,\epsilon}$ are probability densities:
$\left(1-\epsilon\right)\left\{P_{0}\left[p_{1}/p_{0}<c^{\prime\prime}\right]+\frac{1}{c^{\prime\prime}}P_{1}\left[p_{1}/p_{0}\geq c^{\prime\prime}\right]\right\}$ $=1,$ (16)
$\left(1-\epsilon\right)\left\{P_{1}\left[p_{1}/p_{0}>c^{\prime}\right]+c^{\prime}P_{0}\left[p_{1}/p_{0}\leqslant c^{\prime}\right]\right\}$ $=1.$ (17)
We assume that $\epsilon$ is small enough so that $H^{\epsilon}_{0}\cap H^{\epsilon}_{1}=\emptyset.$ Then, the above ensures $c^{\prime}<c^{\prime\prime}$ and $\frac{q_{1,\epsilon}(x)}{q_{0,\epsilon}(x)}=\max\{c^{\prime},\min\{c^{\prime\prime},q_{1,\epsilon}(x)/q_{0,\epsilon}(x)\}\}$. Huber used $R_{n,\epsilon}=\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$ for deriving risk optimal tests. Then we have the following as a corollary of Theorem 1 of [1] and our Theorem 2.1.
Corollary 2.4.
$\{R_{n,\epsilon}\}_{n}$ is a test supermartingale for the adaptive null contamination model $H^{\epsilon,\infty}_{0}$. Hence, recalling (11), the stopping time
$\tau^{*}=\inf\{n:R_{n,\epsilon}\geq 1/\alpha\}$ (18)
at which we reject the null yields a level-$\alpha$ sequential test. Moreover, it achieves the optimal growth rate under the adaptive alternative contamination model $H^{\epsilon,\infty}_{1},$ i.e.,
$\inf_{\mathbf{P}\in H^{\epsilon,\infty}_{1}}\mathbb{E}_{\mathbf{P}}\!\left(\log(R_{n,\epsilon})/{n}\right)=\inf_{{P}\in H^{\epsilon}_{1}}\mathbb{E}_{P}\!\left(\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right)=D_{\operatorname{KL}}(Q_{1,\epsilon},Q_{1,\epsilon}).$ (19)
As discussed in [37, Chapter 10], LFDs can be constructed for a broader class of models, of which (1) and (2) arise as special cases. Consequently, the above result also extends to other types of robust models—such as those defined via $\epsilon$-neighborhoods around two distributions $P_{0}$ and $P_{1}$ in Kolmogorov, Lévy, Prohorov metrics, etc., when $\frac{p_{1}(x)}{p_{0}(x)}$ is monotone in $x$.
3 Robust tests for composite nulls and alternatives when an LFD pair exists
In this section, we consider both the null $\mathcal{P}_{0}$ and the alternative $\mathcal{P}_{1}$ to be composite and assume that an LFD pair exists for testing between these two hypotheses. To handle small deviations from the idealized models, we test $\mathcal{P}^{\epsilon}_{0}$ vs. $\mathcal{P}^{\epsilon}_{1}$, where for $j=0,1$,
$\mathcal{P}^{\epsilon}_{j}=\bigcup_{P\in\mathcal{P}_{j}}B({P},{\epsilon}),\penalty 10000\ \penalty 10000\$$B({P},{\epsilon})=\{Q\in\mathcal{M}:Q=(1-\epsilon)P+\epsilon H,H\in\mathcal{M}\}\text{ or }$ (20)
$B({P},{\epsilon})=\{Q\in\mathcal{M}:D_{\operatorname{TV}}(P,Q)\leqslant\epsilon\}.$ (21)
Theorem 3.1.
Let $(P_{0}^{*},P_{1}^{*})$ be the least favorable pair in terms of risk defined in (4) for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ and $(Q_{0,\epsilon},Q_{1,\epsilon})$ be the least favorable pair in terms of risk for testing $B({P}_{0}^{*},{\epsilon})$ vs. $B({P}_{1}^{*},{\epsilon})$, which are $\epsilon$ neighborhoods around $P_{0}^{*}$ and $P_{1}^{*}$ respectively, as defined in (20) or (21). Then, for any fixed sample size $n$, $(Q_{0,\epsilon},Q_{1,\epsilon})$ is the least favorable pair in terms of risk for testing $\mathcal{P}_{0}^{\epsilon}$ vs. $\mathcal{P}_{1}^{\epsilon}$.
Proof.
Since $(P_{0}^{*},P_{1}^{*})$ is the least favorable pair for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ we have, $R(P_{0}^{*},\phi)\geq R(P,\phi),$ $\forall P\in\mathcal{P}_{0},$ and for all likelihood ratio test $\phi$ between ${P}_{0}^{*}$ and ${P}_{1}^{*}$, which implies
$P_{0}^{*}\left[\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\geq\eta\right]\geq P\left[\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\geq\eta\right]\penalty 10000\ \forall\eta\in\mathbb{R},\forall P\in\mathcal{P}_{0}.$ (22)
From [1], we know that $\frac{q_{1,\epsilon}}{q_{0,\epsilon}}=\max\{c_{1},\min\{c_{2},\frac{p_{1}^{*}}{p_{0}^{*}}\}\}$, for some $c_{1}\leqslant c_{2}$. So, for $c_{1}\leqslant\eta\leqslant c_{2},$
$P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]=P\left[\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\geq\eta\right]\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]=P_{0}^{*}\left[\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\geq\eta\right]$ (23)
Hence, for any $c_{1}\leqslant\eta\leqslant c_{2},$
$P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\geq P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\penalty 10000\ \forall P\in\mathcal{P}_{0}.$ (24)
Also, note that $P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]=0=P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right],$ for any $\eta>c_{2}$ and $P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]=1=P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right],$ for any $\eta<c_{1}$. Therefore, (24) holds for any $\eta\in\mathbb{R}$.
Now, for any $Q\in\mathcal{P}_{0}^{\epsilon}$, there exists $P^{\prime}\in\mathcal{P}_{0}$ such that $Q\in B(P^{\prime},\epsilon)$.
For the contamination model (20), for any $\eta\in\mathbb{R},$
$Q_{0,\epsilon}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$ $=\max_{P\in B(P_{0}^{*},\epsilon)}\penalty 10000\ P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$
$=\max_{H}\left[(1-\epsilon)P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]+\epsilon H\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\right]$
$=(1-\epsilon)P_{0}^{*}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]+\epsilon$
$\geq(1-\epsilon)P^{\prime}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]+\epsilon$
$=\max_{P\in B(P^{\prime},\epsilon)}\penalty 10000\ P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$
$\geq Q\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right].$
Similarly, for the TV model (21), for any $\eta\in\mathbb{R},$
$Q_{0,\epsilon}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$ $=\sup_{P\in B(P_{0}^{*},\epsilon)}\penalty 10000\ P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$
$=P^{*}_{0}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]+\epsilon\quad(\text{By definition of TV distance})$
$\geq P^{\prime}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]+\epsilon$
$=\sup_{P\in B(P^{\prime},\epsilon)}\penalty 10000\ P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]$
$\geq Q\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right].$
Therefore, for both cases, we have $Q_{0,\epsilon}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\geq Q\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right],\forall\eta\in\mathbb{R},\forall Q\in\mathcal{P}_{0}^{\epsilon}$ and similarly, one can show that for both cases, $Q_{0,\epsilon}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}>\eta\right]\geq Q\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}>\eta\right],\forall\eta\in\mathbb{R},\forall Q\in\mathcal{P}_{0}^{\epsilon}$.
Suppose, we have i.i.d. samples $X_{1},\cdots,X_{n}$, for any fixed $n$.
Let $F$ and $G$ be the CDFs of $\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$, under $Q_{0,\epsilon}$ and $Q$ respectively. From the stochastic monotonicity established above, we have $F\leqslant G$. Now, let $Z_{1},\cdots,Z_{n}$ are i.i.d from Unif(0,1)(0,1). Then, $F^{-1}(Z_{i})\geq G^{-1}(Z_{i})$ and $F^{-1}(Z_{i})\overset{\text{d}}{=}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$ when $X_{i}\sim Q_{0,\epsilon}$ and $G^{-1}(Z_{i})\overset{\text{d}}{=}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$, when $X_{i}\sim Q$, $i=1,\cdots,n$. Therefore, we obtain
$Q_{0,\epsilon}\left[\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}>\eta\right]\geq Q\left[\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}>\eta\right],\forall\eta\in\mathbb{R},\forall Q\in\mathcal{P}_{0}^{\epsilon},$ (25)
and
$Q_{0,\epsilon}\left[\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq\eta\right]\geq Q\left[\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq\eta\right],\forall\eta\in\mathbb{R},\forall Q\in\mathcal{P}_{0}^{\epsilon}$ (26)
Note that for any likelhihood ratio $\phi^{*}$ between $Q_{0,\epsilon}$ and $Q_{1,\epsilon}$, there exists $\eta\in\mathbb{R},\gamma\in[0,1]$ such that
$\phi^{*}(x_{1},\cdots,x_{n})=\begin{cases}1&\text{if }\prod_{i=1}^{n}\frac{q_{1,\epsilon}}{q_{0,\epsilon}}(x_{i})>\eta,\\ \gamma&\text{if }\prod_{i=1}^{n}\frac{q_{1,\epsilon}}{q_{0,\epsilon}}(x_{i})=\eta,\\ 0&\text{if }\prod_{i=1}^{n}\frac{q_{1,\epsilon}}{q_{0,\epsilon}}(x_{i})<\eta.\end{cases}$ (27)
Then,
$R(Q_{0,\epsilon},\phi^{*})$ $=C_{0}\times\left[Q_{0,\epsilon}\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}>\eta\right)+\gamma\times Q_{0,\epsilon}\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}=\eta\right)\right]$
$=C_{0}\times\left[(1-\gamma)\times Q_{0,\epsilon}\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}>\eta\right)+\gamma\times Q_{0,\epsilon}\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq\eta\right)\right]$
$\geq C_{0}\times\left[(1-\gamma)\times Q\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}>\eta\right)+\gamma\times Q\left(\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq\eta\right)\right]$
$=R(Q,\phi^{*}),\forall Q\in\mathcal{P}_{0}^{\epsilon}$
The proof for the part $R(Q_{1,\epsilon},\phi^{*})\geq R(Q,\phi^{*})$ is identical. ∎
Remark 3.2.
The above theorem holds for sequential probability ratio tests as well We first consider Wald’s SPRT, $\phi^{*}_{\text{SPRT}}$ with stooping rule $\tau=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\notin(A,B)\}$. Define, $\tau_{1}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq B\}$ and $\tau_{0}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\leqslant A\}$. Let $F$ and $G$ be the CDFs of $\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$, under $Q_{0,\epsilon}$ and $Q$ respectively. For $Z_{1},\cdots,Z_{n}\overset{i.i.d.}{\sim}U(0,1)$, define $\tau^{\prime}_{1}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}F^{-1}(Z_{i})\geq B\}$ and $\tau^{\prime\prime}_{1}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}G^{-1}(Z_{i})\geq B\}$. Also, $\tau^{\prime}_{0}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}F^{-1}(Z_{i})\leqslant A\}$ and $\tau^{\prime\prime}_{0}=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}G^{-1}(Z_{i})\leqslant A\}$. Then, it follows from the stochastic monotonicity established in the proof of the theorem that $F\leqslant G$ and hence, $\tau^{\prime}_{0}\geq\tau^{\prime\prime}_{0}$ and $\tau^{\prime}_{1}\leqslant\tau^{\prime\prime}_{1}$. So,
$R(Q_{0,\epsilon},\phi^{*}_{\text{SPRT}})$ $=C_{0}Q_{0,\epsilon}(\phi^{*}_{\text{SPRT}}\text{ accepts }H_{1})$
$=C_{0}Q_{0,\epsilon}\left(\tau_{1}<\tau_{0}\right)$
$=C_{0}\mathbb{P}_{Z_{i}\overset{i.i.d.}{\sim}U(0,1)}\left(\tau_{1}^{\prime}<\tau_{0}^{\prime}\right)$
$\geq C_{0}\mathbb{P}_{Z_{i}\overset{i.i.d.}{\sim}U(0,1)}\left(\tau_{1}^{\prime\prime}<\tau_{0}^{\prime}\right)$
$\geq C_{0}\mathbb{P}_{Z_{i}\overset{i.i.d.}{\sim}U(0,1)}\left(\tau_{1}^{\prime\prime}<\tau_{0}^{\prime\prime}\right)$
$=C_{0}Q(\tau_{1}<\tau_{0})$
$=R(Q,\phi^{*}_{\text{SPRT}}),$
for all $Q\in\mathcal{P}_{0}^{\epsilon}$. Simliarly, one can show $R(Q_{1,\epsilon},\phi^{*}_{\text{SPRT}})\geq R(Q,\phi^{*}_{\text{SPRT}})$, for all $Q\in\mathcal{P}_{1}^{\epsilon}$.
Similarly, for the one-sided SPRT, $\phi^{*}_{\text{S}}$ with $\tau=\inf\{n\in\mathbb{N}:\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}\geq B\}$, we show the same:
$R(Q_{0,\epsilon},\phi^{*}_{\text{S}})=C_{0}Q_{0,\epsilon}\left(\tau<\infty\right)=C_{0}\mathbb{P}_{Z_{i}\overset{iid.}{\sim}U(0,1)}\left(\tau^{\prime}_{1}<\infty\right)\geq C_{0}\mathbb{P}_{Z_{i}\overset{iid.}{\sim}U(0,1)}\left(\tau^{\prime\prime}_{1}<\infty\right)=R(Q,\phi^{*}_{\text{S}}),$
for all $Q\in\mathcal{P}_{0}^{\epsilon}$.
Remark 3.3.
Note that for the above theorem to hold, it is not necessary for the neighborhoods $B(P,\epsilon)$ to be $\epsilon$-contamination or total variation (TV) balls. The only properties of $B(\cdot,\epsilon)$ required in the proof are the following: an LFD pair exists for testing $B(\cdot,\epsilon)$ vs. $B(\cdot,\epsilon)$ as long as they do not intersect each other and for any $Q,Q^{\prime}\in\mathcal{P}_{0}\cup\mathcal{P}_{1},\eta\in\mathbb{R}$ satisfying
$Q\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\geq Q^{\prime}\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\implies\displaystyle\sup_{P\in B(Q,\epsilon)}P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right]\geq\sup_{P\in B(Q^{\prime},\epsilon)}P\left[\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\geq\eta\right].$
Clearly, both $\epsilon$-contamination and TV neighborhoods satisfy this condition, as shown in the proof. However, it is worth noting that other robust models, eg., $\epsilon$-balls based on Kolmogorov or Lévy distances, also satisfy it, under the assumption that $\frac{p_{1}(x)}{p_{0}(x)}$ is monotone in $x$, for all $P_{0}\in\mathcal{P}_{0},P_{1}\in\mathcal{P}_{1}$.
3.1 Adaptive contamination models
Recalling the adaptive contamination model for the null from Section 2.1, we define the adaptive contamination model for the composite hypothesis, $\mathcal{H}^{\epsilon,\infty}_{i}$ to be the set of all possible joint distribution of the sequence $X_{1},X_{2},\cdots$ such that the conditional distribution of $X_{n}$ given $X^{n-1}$ lies within $\mathcal{H}^{\epsilon}_{i}$. So, at each $n$, the adversary defines a measurable mapping $Q_{n}:\Omega^{n-1}\to\mathcal{H}^{\epsilon}_{i}$ and $X_{n}$ is drawn from the distribution $Q_{n}(x^{n-1}),$ given that $X^{n-1}=x^{n-1}$.
Our test retains its validity properties under the null hypothesis and growth rate under the alternative, even in the presence of an adaptive contamination model, where $H^{\epsilon}_{i}$ is some $\epsilon$ neighborhood around the hypothesized distribution $P_{i}$, for example, $\epsilon$-contamination or total variation balls defined in (20) or (21).
3.2 Robust test for composite null vs. composite alternative
Let $(P_{0},P_{1})$ be the least favorable pair for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ and $(Q_{0,\epsilon},Q_{1,\epsilon})$ be the least favorable pair in terms of risk for testing $B({P}_{0},{\epsilon})$ vs. $B({P}_{1},{\epsilon})$, defined in (20). Then, with $R_{n,\epsilon}=\prod_{i=1}^{n}\frac{q_{1,\epsilon}(X_{i})}{q_{0,\epsilon}(X_{i})}$, we have the following as a corollary of Theorem 1 of [1] and our Theorems 2.1 and 3.1.
Corollary 3.4.
$\{R_{n,\epsilon}\}_{n}$ is a test supermartingale for the adaptive null contamination model $H^{\epsilon,\infty}_{0}$. Hence, recalling (11), the stopping time
$\tau^{*}=\inf\{n:R_{n,\epsilon}\geq 1/\alpha\}$ (28)
at which we reject the null yields a level-$\alpha$ sequential test. Moreover it achieve the optimal growth rate, i.e.,
$\inf_{\mathbf{P}\in H^{\epsilon,\infty}_{1}}\mathbb{E}_{\mathbf{P}}\!\left(\log(R_{n,\epsilon})/{n}\right)=\inf_{{P}\in H^{\epsilon}_{1}}\mathbb{E}_{P}\!\left(\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right)=D_{\operatorname{KL}}(Q_{1,\epsilon},Q_{1,\epsilon}).$
As discussed in Remark 3.3, similar results can also be established for other robust models, e.g., $\epsilon$-balls based on total variation, Kolmogorov or Levy distance.
4 Robust tests for composite nulls and alternatives when an LFD pair does not exist
Now, we consider the problem of robust hypothesis testing, where an LFD pair does not exist for the pre-contamination composite null and composite alternative.
4.1 Robust predictable plug-in for composite alternatives
In this section, we address the challenge of robustly testing the composite alternatives. Let the idealized models be $P_{0}$ vs. $\mathcal{P}_{1}$, where $\mathcal{P}_{1}$ is a set of distribution functions that does not include $P_{0}$. We will assume that there exists a common reference measure $\mu$ for $\mathcal{P}_{1}$ and $P_{0}$ so that we can associate the distributions with their densities.
To handle a composite alternative hypothesis, one natural way is to attempt to learn it from past observations, at each round $n$, and plug it in as an estimate of the alternative distribution based on past observations. This is known as the “predictable plug-in” or simply “plug-in” method. This method is originally due to [38], which has recently been used for handling various parametric and non-parametric composite alternatives [39, 21, 40].
However, we often encounter deviations from these idealized models in real-world data, potentially due to contamination or deviations from the assumed distribution. To navigate this problem, we use a robust estimate of the alternative distribution, which lies in $\mathcal{P}_{1}$, where the actual data might come from $\epsilon$-neighbourhood of some distribution in the alternative $\mathcal{P}_{1}$, which is
[…]
$\mathcal{P}^{\epsilon}_{1}=\bigcup_{P_{1}\in\mathcal{P}_{1}}\{Q\in\mathcal{M}:D_{\operatorname{TV}}(P_{1},Q)\leqslant\epsilon\}.$ (30)
4.1.1 An adaptive contamination model for the null
Under the null, we assume that the data $X_{1},X_{2},\dots$ is generated from some distribution in $H^{\epsilon,\infty}_{0}$ defined in Section 2.1, i.e., while $P_{0}$ is fixed and known, the unknown contaminations may vary across time. In fact, the contamination at time $n$ may depend on the previously observed data, as noted earlier.
Let $\hat{p}_{1,n}$ be some robust estimate of the density of the data based on past observations $X^{n-1}$, which belongs to the alternative. Let $\hat{P}_{1,n}$ be the distribution corresponding to the density $\hat{p}_{1,n}$, with $\hat{P}_{1,n}\in\mathcal{P}_{1}$.
Now we define $\hat{q}_{n,j,\epsilon}$ similarly as ${q}_{j,\epsilon}$ was defined in (15) (for $j=0,1$) .
[…]
We calculate the numbers $0\leqslant c^{\prime}_{n},c^{\prime\prime}_{n}\leqslant\infty$, that are determined such that $\hat{q}_{n,0,\epsilon},\hat{q}_{n,1,\epsilon}$ are probability densities:
$\left(1-\epsilon\right)\left\{P_{0}\left[\hat{p}_{1,n}/p_{0}<c^{\prime\prime}_{n}\right]+\frac{1}{c^{\prime\prime}_{n}}\hat{P}_{1,n}\left[\hat{p}_{1,n}/p_{0}\geq c^{\prime\prime}_{n}\right]\right\}$ $=1,$ (33)
$\left(1-\epsilon\right)\left\{\hat{P}_{1,n}\left[\hat{p}_{1,n}/p_{0}>c^{\prime}_{n}\right]+c^{\prime}_{n}P_{0}\left[\hat{p}_{1,n}/p_{0}\leqslant c^{\prime}_{n}\right]\right\}$ $=1.$ (34)
Note that if $c^{\prime}_{n}<c^{\prime\prime}_{n}$, we have ${\frac{\hat{q}_{n,1,\epsilon}(x)}{\hat{q}_{n,0,\epsilon}(x)}}=\max\{c^{\prime}_{n},\min\{c^{\prime\prime}_{n},\frac{\hat{p}_{1,n}(x)}{p_{0}(x)}\}\}$. We now define $R_{0,\epsilon}^{\text{plug-in}}=1$ and $R_{n,\epsilon}^{\text{plug-in}}=R_{n-1,\epsilon}^{\text{plug-in}}\times\hat{E}_{n,\epsilon}(X_{n}),n=1,2,\dots$ where
[…]
In other words, $R_{n,\epsilon}^{\text{plug-in}}$ is the product of $\hat{E}_{n,\epsilon}$. Thus, if each $\hat{E}_{n,\epsilon}$ is an e-variable (conditioned on the past), $R_{n,\epsilon}^{\text{plug-in}}$ will be a test supermartingale. The following short calculation shows that this is indeed the case:
$\mathbb{E}_{Q_{n}}[\hat{E}_{n,\epsilon}(X_{n})\mid X^{n-1}=x^{n-1}]$
$=\mathbbm{I}_{c^{\prime}_{n}\geq c^{\prime\prime}_{n}}+\frac{\mathbb{E}_{Q_{n}}\left[\frac{\hat{q}_{n,1,\epsilon}(X_{n})}{\hat{q}_{n,0,\epsilon}(X_{n})}\mid X^{n-1}=x^{n-1}\right]}{\mathbb{E}_{P_{0}}\left[\frac{\hat{q}_{n,1,\epsilon}(X)}{\hat{q}_{n,0,\epsilon}(X)}\mid X^{n-1}=x^{n-1}\right]+(c^{\prime\prime}_{n}-c^{\prime}_{n})\epsilon}\mathbbm{I}_{c^{\prime}_{n}<c^{\prime\prime}_{n}}$
$\leqslant\mathbbm{I}_{c^{\prime}_{n}\geq c^{\prime\prime}_{n}}+\mathbbm{I}_{c^{\prime}_{n}<c^{\prime\prime}_{n}}=1.$
The last inequality follows from the fact that the total variation distance is an integral probability metric in the sense that for any pair of real numbers $c_{1}<c_{2}$,
$D_{\operatorname{TV}}(P,Q)=\frac{1}{c_{2}-c_{1}}\sup_{c_{1}\leq f\leq c_{2}}\left|\operatorname*{\mathbb{E}}_{X\sim P}f(X)-\operatorname*{\mathbb{E}}_{X\sim Q}f(X)\right|,$ (36)
since the conditional distribution $Q_{n}=Q_{n}(x^{n-1})$ of $X_{n}$ conditioned on $X^{n-1}=x^{n-1}$ satisfies $Q_{n}\in H^{\epsilon}_{0},\text{ i.e., }D_{\operatorname{TV}}(Q_{n},P_{0})\leqslant\epsilon$. We immediately conclude the following.
Theorem 4.1.
$R_{n,\epsilon}^{\text{plug-in}}$ is a test supermartingale for the adaptive null contamination model $H^{\epsilon,\infty}_{0}$. Hence, recalling (11), the stopping time
$\tau^{*}=\inf\{n:R_{n,\epsilon}^{\text{plug-in}}\geq 1/\alpha\}$ (37)
at which we reject the null yields a level-$\alpha$ sequential test.
4.1.2 Growth rate analysis
For analyzing the growth rate of this test supermartingale, we assume that the estimator $\hat{p}_{1,n}$ converges pointwise in the following sense:
Assumption 4.2.
If $X_{1},X_{2},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}H$, for some distribution $H\in\mathcal{P}^{\epsilon}_{1}$, then $\hat{p}_{1,n}\to p_{1}^{H}$ almost surely as $n\to\infty$, where $p_{1}^{H}$ is a density function (which may depend on $H$) corresponding to some distribution $P_{1}^{H}\in\mathcal{P}_{1}$.
Note that the above pointwise convergence assumption is much weaker than assuming the existence of a pointwise consistent estimator since $P_{1}^{H}$ is any arbitrary distribution in $\mathcal{P}_{1}$. Pointwise convergence holds in a wide variety of estimation problems within the robust statistics literature. For instance, robust M-estimators for certain parametric testing problems are known to be strongly consistent under regularity conditions [37, Chapter 6]. To illustrate with a concrete example, when estimating the mean of a Gaussian distribution under contamination, the sample median serves as a robust estimator that converges almost surely to the population median.
The next theorem shows that $\log R_{n,\epsilon}^{\text{plug-in}}/n$ converges almost surely, under some assumptions.
Theorem 4.3.
Suppose that $X_{1},X_{2},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}H$, for some distribution $H\in\mathcal{P}^{\epsilon}_{1}$ (where $\mathcal{P}^{\epsilon}_{1}$ is as defined in (29) or (30)) and the estimator $\hat{p}_{1,n}$ satisfies Assumption 4.2. We further assume that $P(p^{H}_{1}/p_{0}=c)=0$, for all $c\in\mathbb{R}$ and for $P\in P_{0}\cup\mathcal{P}_{1}$. We consider $\epsilon$ being sufficiently small so that $c^{\prime\prime}_{H}>c^{\prime}_{H}$, where $c^{\prime\prime}_{H},c^{\prime}_{H}$ are solutions of (41) and (42), respectively, with $P_{1}$ being replaced by $P_{1}^{H}$. Then, as $n\to\infty$,
$\frac{1}{n}\log R_{n,\epsilon}^{\text{plug-in}}\to r_{H,\epsilon}^{\text{plug-in}}\text{ almost surely,}$
where $r_{H,\epsilon}^{\text{plug-in}}\geq D_{\operatorname{KL}}(Q^{H}_{1,\epsilon},Q^{H}_{0,\epsilon})-2(\log c^{\prime\prime}_{H}-\log c^{\prime}_{H})\epsilon-\log(1+2(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon)$ and $Q_{1,\epsilon}^{H},Q_{0,\epsilon}^{H}$ are the distributions with densities $q_{1,\epsilon}^{H},q_{0,\epsilon}^{H}$ as defined in (14) and (15) with $P_{1}$ being replaced by $P_{1}^{H}$.
Remark 4.4.
In Theorem 4.3, if we further assume that the maximum bias $b_{P_{1}}(\epsilon,x):=\sup_{H:D_{\operatorname{TV}}(H,P_{1})\leqslant\epsilon}|p_{1}^{H}(x)-p_{1}(x)|$ is a real-valued function such that $\lim_{\epsilon\to 0}b_{P_{1}}(\epsilon,x)=0$, for all $x\in\mathbb{R}$, we still could not conclude in general that $\inf_{H:D_{\operatorname{TV}}(H,P_{1})\leqslant\epsilon}r_{H,\epsilon}^{\text{plug-in}}$ converges to the oracle (when $P_{1}$ is known) growth rate $D_{\operatorname{KL}}(P_{1},P_{0})$, because of the difficulty in interchanging the expectation and the limit $\epsilon\to 0$. However, under the aforementioned vanishing bias assumption, in specific situations, such as exponential family distributions, one can show that
$\lim_{\epsilon\to 0}\inf_{H:D_{\operatorname{TV}}(H,P_{1})\leqslant\epsilon}r_{H,\epsilon}^{\text{plug-in}}=D_{\operatorname{KL}}(P_{1},P_{0}).$
4.2 Robust numeraire: composite null vs. simple alternative
In this section, we consider the null $\mathcal{P}_{0}$ to be composite and the alternative to be simple $P_{1}$. To account for the small deviations from the idealized models, consider the following:
$\mathcal{P}^{\epsilon}_{0}=\bigcup_{P\in\mathcal{P}_{0}}\{Q\in\mathcal{M}:Q=(1-\epsilon)P+\epsilon H,H\in\mathcal{M}\},\quad$ (38)
 $\text{ or }\mathcal{P}^{\epsilon}_{0}=\bigcup_{P\in\mathcal{P}_{0}}\{Q\in\mathcal{M}:D_{\operatorname{TV}}(P,Q)\leqslant\epsilon\},$ (39)
and ${H}^{\epsilon}_{1}$ is either (1) or (2). Recalling the adaptive contamination model for the null from Section 2.1, we define the adaptive contamination model for the composite null, $\mathcal{H}^{\epsilon,\infty}_{0}$ to be the set of all possible joint distribution of the sequence $X_{1},X_{2},\cdots$ such that the conditional distribution of $X_{n}$ given $X^{n-1}$ lies within $\mathcal{H}^{\epsilon}_{0}$. So, at each $n$, the adversary defines a measurable mapping $Q_{n}:\Omega^{n-1}\to\mathcal{P}^{\epsilon}_{0}$ and $X_{n}$ is drawn from the distribution $Q_{n}(x^{n-1}),$ given that $X^{n-1}=x^{n-1}$.
4.2.1 The reverse information projection (RIPr) and numeraire
For testing composite nulls versus a simple alternative, the reverse information projection (RIPr) has recently emerged as an central tool [2, 6, 4]. In his PhD thesis, Li [5] defined the RIPr as follows: if we assume that $\mathcal{P}_{0}$ is convex and has a reference measure, and $\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{1},P)<\infty$, then there exists a unique sub-probability measure $P_{0}$ that satisfies
$D_{\operatorname{KL}}(P_{1},P_{0})=\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{1},P).$ (40)
This $P_{0}$ is defined as the RIPr of $P_{1}$ on $\mathcal{P}_{0}$.
Recently, [4] significantly generalized this definition by exploiting a certain duality of the RIPr with e-variables. They first show that for any null $\mathcal{P}_{0}$ and simple alternative $P_{1}$, there always exists a ($P_{1}$ almost surely) unique and strictly positive e-variable $B^{*}$ called the numeraire, such that for any other e-variable $B$ for $\mathcal{P}_{0}$, $\mathbb{E}_{P_{1}}[B/B^{*}]\leqslant 1$ (“numeraire” property, closely related to Cover’s optimal portfolios). An equivalent statement is that for any e-variable $B$ for $\mathcal{P}_{0}$, $\mathbb{E}_{P_{1}}\log(B^{*}/B)\geq 0$ (“log-optimality”; see also [2, 6]). This implies that the growth rate of $B^{*}$, $\mathbb{E}_{P_{1}}\log(B^{*})$, is larger than that of any other e-variable $B$. It is somewhat remarkable that the existence and uniqueness of such an optimal e-variable can be established under absolutely no assumptions on $\mathcal{P}_{0}$ and $P_{1}$. [4] then define a measure $P_{0}$ by defining its likelihood ratio (Radon-Nikodym derivative) with respect to $P_{1}$, $\frac{dP_{0}}{dP_{1}}:=\frac{1}{B^{*}}.$ This $P_{0}$ is the more general definition of the Reverse Information Projection (RIPr) of $P_{1}$ onto $\mathcal{P}_{0}$, and it can be shown to match Li’s original definition under his additional assumptions (of convexity, reference measure, and finiteness of $D_{\operatorname{KL}}$). Further, a very general strong duality holds between $B^{*}$ and $P_{0}$ that we omit here for brevity.
In what follows, the RIPr $P_{0}$ is the sole representative of the composite null $\mathcal{P}_{0}$, and even though it may be a sub-probability measure in general, we can still proceed as if we were dealing with a simple null. We elaborate below.
4.2.2 Robustifying the numeraire
Let ${P}_{0}$ be the RIPr of ${P}_{1}$ on the null $\mathcal{P}_{0}$. Suppose, for $j=1,2$, $p_{j}$ be the density of $P_{j}$ with respect to some common dominating measure $\mu$. Let, $k=\int p_{0}d\mu$. Note that since ${P}_{0}$ is a sub-probability measure, we have $k\leqslant 1$. We obtain ${q}_{j,\epsilon}$ as defined in (14) and (15). We calculate the numbers $0\leqslant c^{\prime},c^{\prime\prime}\leqslant\infty$ are determined such that $q_{0,\epsilon}$ is a sub-probability density with $\int q_{0,\epsilon}d\mu=k$ and $q_{1,\epsilon}$ is a probability density:
$\left(1-\epsilon\right)\left\{P_{0}\left[p_{1}/p_{0}<c^{\prime\prime}\right]+\frac{1}{c^{\prime\prime}}P_{1}\left[p_{1}/p_{0}\geq c^{\prime\prime}\right]\right\}$ $=k,$ (41)
$\left(1-\epsilon\right)\left\{P_{1}\left[p_{1}/p_{0}>c^{\prime}\right]+c^{\prime}P_{0}\left[p_{1}/p_{0}\leqslant c^{\prime}\right]\right\}$ $=1.$ (42)
We choose $\epsilon$ sufficiently small, then $c^{\prime}<c^{\prime\prime}$ and we have ${\frac{{q}_{1,\epsilon}(x)}{q_{0,\epsilon}(x)}}=\max\{c^{\prime},\min\{c^{\prime\prime},\frac{{p}_{1}(x)}{p_{0}(x)}\}\}$. Define $R_{0,\epsilon}^{\text{RIPr}}=1$ and $R_{n,\epsilon}^{\text{RIPr}}=R_{n-1,\epsilon}^{\text{RIPr}}\times B_{\epsilon}(X_{n}),n=1,2,\dots$ where
$B_{\epsilon}(x):=\frac{\frac{{q}_{1,\epsilon}(x)}{q_{0,\epsilon}(x)}}{\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]+(c^{\prime\prime}-c^{\prime})\epsilon}.$ (43)
The term $\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]$ might not have a closed form expression, in that case, we need to rely on numerical approximations in practice. In parametric setting, one can discretize the parameter space in $\mathcal{P}_{0}$ by selecting a grid of finitely many representative points, and at each grid point, we approximate the expectation using Monte Carlo sampling, where we generate independent samples from the corresponding distribution and compute the empirical mean of the likelihood ratio. The final approximation is then obtained by taking the maximum over all the values at the grid points. However, in many simple settings, we have that the RIPr satisfies the maximum, i.e.,
$\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]=\mathbb{E}_{X\sim P_{0}}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right].$ (44)
When this condition holds, computing the test supermartingale becomes significantly simpler. Next, we provide a sufficient condition under which this equality is guaranteed.
Proposition 4.5.
Suppose that $\{P_{\theta}:\theta\in\Theta\}$ is some exponential family distribution with one-dimensional sufficient statistic for the parameter $\theta$. Let ${P}_{1}=P_{\theta_{1}}$ and $\mathcal{P}_{0}=\{P_{\theta}:\theta\in\Theta_{0}\}$ are such that $\sup\Theta_{0}\leqslant\theta_{1}$ (or $\inf\Theta_{0}\geq\theta_{1}$). Then, the RIPr, $P_{0}$, is $P_{\theta_{0}^{*}}$, where $\theta_{0}^{*}=\sup\Theta_{0}$ (or $\theta_{0}^{*}=\inf\Theta_{0}$), and (44) holds.
The above result also holds for a general monotone likelihood ratio family, eg., see [41] [Chapter 2.6.1].
Now, we observe that, under our adaptive null contamination model, for any possible conditional distribution of $X_{n}$ conditioned on $X^{n-1}=x^{n-1}$ (denoted as $Q_{n}=Q_{n}(x^{n-1})$), there exists $P_{0,n}\in\mathcal{P}_{0}$, such that $D_{\operatorname{TV}}(Q_{n},P_{0,n})\leqslant\epsilon$. Then,
$\mathbb{E}_{Q_{n}}[B_{\epsilon}(X_{n})\mid X^{n-1}=x^{n-1}]$
$=\mathbb{E}_{Q_{n}}\left[\frac{\frac{{q}_{1,\epsilon}(X_{n})}{q_{0,\epsilon}(X_{n})}}{\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]+(c^{\prime\prime}-c^{\prime})\epsilon}\mid X^{n-1}=x^{n-1}\right]$
$\leqslant\frac{\mathbb{E}_{Q_{n}}\left[\frac{{q}_{1,\epsilon}(X_{n})}{q_{0,\epsilon}(X_{n})}\mid X^{n-1}=x^{n-1}\right]}{\mathbb{E}_{P_{0,n}}\left[\frac{{q}_{1,\epsilon}(X_{n})}{q_{0,\epsilon}(X_{n})}\right]+(c^{\prime\prime}-c^{\prime})\epsilon}$
$\leqslant 1.$
The last inequality follows from (36), since the conditional distribution $Q_{n}=Q_{n}(x^{n-1})$ satisfies $D_{\operatorname{TV}}(Q_{n},P_{0,n})\leqslant\epsilon$. We immediately conclude the following.
Theorem 4.6.
Suppose that $\epsilon>0$ is sufficiently small such that we have $c^{\prime\prime}>c^{\prime}$. Then, $R_{n,\epsilon}^{\text{RIPr}}$ is a test supermartingale for the adaptive null contamination model $\mathcal{H}^{\epsilon,\infty}_{0}$. Hence, recalling (11), the stopping time
$\tau^{*}=\inf\{n:R_{n,\epsilon}^{\text{RIPr}}\geq 1/\alpha\}$ (45)
at which we reject $\mathcal{P}^{\epsilon}_{0}$ yields a level-$\alpha$ sequential test.
It turns out that similar results can be established for the composite null case, akin to those observed in the simple null versus simple alternative scenario.
Theorem 4.7.
Suppose that $\epsilon>0$ be sufficiently small such that we have $c^{\prime\prime}>c^{\prime}$ and $X_{1},X_{2},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}Q\in H^{\epsilon}_{1}$. Then, as $n\to\infty$,
$\frac{\log R_{n,\epsilon}^{\text{RIPr}}}{n}\to r^{Q,\epsilon}_{\text{RIPr}}\text{ almost surely for some constant }r^{Q,\epsilon}_{\text{RIPr}}$
and the growth rate,
$r^{\epsilon}_{\text{RIPr}}=\inf_{Q\in H^{\epsilon}_{1}}r^{Q,\epsilon}_{\text{RIPr}}\geq D_{\operatorname{KL}}(Q_{1,\epsilon},Q_{0,\epsilon})-2(\log c^{\prime\prime}-\log c^{\prime})\epsilon-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}+(c^{\prime\prime}-c^{\prime})\epsilon\right).$
The next theorem shows that the growth rate of our test converges to $\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{1},P)$, as $\epsilon\to 0$. And it is the optimal e-power or growth rate for testing $\mathcal{P}_{0}$ vs. $P_{1}$ [2, 6].
Theorem 4.8.
If (44) holds, then the growth rate of our test, $r^{\epsilon}_{\text{RIPr}}\to\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{1},P)$, as $\epsilon\to 0$.
This result shows that the growth rate of our robust test supermartingale for composite null is asymptotically optimal as $\epsilon$ approaches zero.
4.3 Combining robust predictable plug-in and robust numeraire: composite null vs. composite alternative
In this section, we address the most general scenario, when both the null $\mathcal{P}_{0}$ and the alternative $\mathcal{P}_{1}$ are composite. We will assume that there exists a common reference measure for $\mathcal{P}_{1}$ and $\mathcal{P}_{0}$ so that we can associate the distributions with their densities. To handle small deviations from the idealized models, we test $\mathcal{P}^{\epsilon}_{0}$ vs. $\mathcal{P}^{\epsilon}_{1}$, defined in (20) or (21).
Our approach combines plug-in and RIPr methods from the last two sections. At each time $n$, let $\hat{p}_{1,n}$ be some robust estimate of the density of the data based on past observations $X^{n-1}$ which belongs to the alternative. Let $\hat{P}_{1,n}$ be the distribution corresponding to the density $\hat{p}_{1,n}$ and $\hat{P}_{1,n}\in\mathcal{P}_{1}$. Let $\hat{P}_{0,n}$ be the reverse information projection (RIPr) of $\hat{P}_{1,n}$ on the null $\mathcal{P}_{0}$.
Now we define $\hat{q}_{n,j,\epsilon}$ similarly as ${q}_{j,\epsilon}$ was defined in (15) (for $j=0,1$) .
[…]
$k_{n}=\int\hat{p}_{0,n}d\mu$. Since ${P}_{0}$ is a sub-probability measure, we have $k_{n}\leqslant 1$. We calculate the numbers $0\leqslant c^{\prime}_{n},c^{\prime\prime}_{n}\leqslant\infty$ are determined such that $\hat{q}_{n,0,\epsilon}$ is a sub-probability density with $\int\hat{q}_{n,0,\epsilon}d\mu=k_{n}$ and $\hat{q}_{n,1,\epsilon}$ is a probability density:
$\hat{P}_{0,n}\left[\hat{p}_{1,n}/\hat{p}_{0,n}<c^{\prime\prime}_{n}\right]+\frac{1}{c^{\prime\prime}_{n}}\hat{P}_{1,n}\left[\hat{p}_{1,n}/\hat{p}_{0,n}\geq c^{\prime\prime}_{n}\right]$ $=\frac{k_{n}}{\left(1-\epsilon\right)},$ (48)
$\hat{P}_{1,n}\left[\hat{p}_{1,n}/\hat{p}_{0,n}>c^{\prime}_{n}\right]+c^{\prime}_{n}\hat{P}_{0,n}\left[\hat{p}_{1,n}/\hat{p}_{0,n}\leqslant c^{\prime}_{n}\right]$ $=\frac{1}{\left(1-\epsilon\right)}.$ (49)
Note that if $c^{\prime}_{n}<c^{\prime\prime}_{n}$, we have ${\frac{\hat{q}_{n,1,\epsilon}(x)}{\hat{q}_{n,0,\epsilon}(x)}}=\max\{c^{\prime}_{n},\min\{c^{\prime\prime}_{n},\frac{\hat{p}_{1,n}(x)}{\hat{p}_{0,n}(x)}\}\}$. Define $R_{0,\epsilon}^{\text{RIPr,plug-in}}=1$ and $R_{n,\epsilon}^{\text{RIPr,plug-in}}=R_{n-1,\epsilon}^{\text{RIPr,plug-in}}\times\hat{B}_{n,\epsilon}(X_{n}),n=1,2,\dots$ where
[…]
As noted in the previous section, $\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{\hat{q}_{n,1,\epsilon}(X)}{\hat{q}_{n,0,\epsilon}(X)}\mid X^{n-1}\right]$ might not have a closed form expression; in that case, we need to rely on numerical approximations. We now show that this construction yields a valid test for the adaptive contamination model $\mathcal{H}^{\epsilon,\infty}_{0}$ defined in the previous section. Under $\mathcal{H}^{\epsilon,\infty}_{0},$ for any possible conditional distribution $Q_{n}=Q_{n}(x^{n-1})$ of $X_{n}$ conditioned on $X^{n-1}=x^{n-1}$, there exists $P_{0,n}\in\mathcal{P}_{0}$, such that $D_{\operatorname{TV}}(Q_{n},P_{0,n})\leqslant\epsilon$. Then,
$\mathbb{E}_{Q_{n}}[\hat{B}_{n,\epsilon}(X_{n})\mid X^{n-1}=x^{n-1}]$
$\leqslant\mathbbm{I}_{c^{\prime}_{n}\geq c^{\prime\prime}_{n}}+\frac{\mathbb{E}_{Q_{n}}\left[\frac{\hat{q}_{n,1,\epsilon}(X_{n})}{\hat{q}_{n,0,\epsilon}(X_{n})}\mid X^{n-1}=x^{n-1}\right]\mathbbm{I}_{c^{\prime}_{n}<c^{\prime\prime}_{n}}}{\mathbb{E}_{P_{0,n}}\left[\frac{\hat{q}_{n,1,\epsilon}(X_{n})}{\hat{q}_{n,0,\epsilon}(X_{n})}\mid X^{n-1}=x^{n-1}\right]+(c^{\prime\prime}_{n}-c^{\prime}_{n})\epsilon}$
$\leqslant\mathbbm{I}_{c^{\prime}_{n}\geq c^{\prime\prime}_{n}}+\mathbbm{I}_{c^{\prime}_{n}<c^{\prime\prime}_{n}}=1.$
The last inequality follows from (36), since the conditional distribution $Q_{n}$ satisfies $D_{\operatorname{TV}}(Q_{n},P_{0,n})\leqslant\epsilon$. We immediately conclude the following.
Theorem 4.9.
$R_{n,\epsilon}^{\text{RIPr,plug-in}}$ is a test supermartingale for the adaptive null contamination model $\mathcal{H}^{\epsilon,\infty}_{0}$. Hence, recalling (11), if
$\tau^{*}=\inf\{n:R_{n,\epsilon}^{\text{RIPr,plug-in}}\geq 1/\alpha\}$ (51)
is the stopping time at which we reject $\mathcal{P}^{\epsilon}_{0}$, then this yields a level-$\alpha$ sequential test.
While the type-I error control holds without any additional assumptions, establishing the growth rate of the test requires additional assumptions. The main challenge lies in the fact that even if we assume $\hat{p}_{1,n}$ to be a strongly consistent estimator, it does not guarantee that the corresponding sequence of reverse information projections, $\hat{p}_{0,n}$, would converge in the almost sure sense. So, we need to explicitly assume the convergence of $\hat{p}_{0,n}$, as formalized below.
Assumption 4.10.
If $\hat{p}_{1,n}$ satisfies Assumption 4.2, then $\hat{p}_{0,n}\to{p}_{0}^{H}$ almost surely as $n\to\infty$, where ${p}_{0}^{H}$ is the density of the measure ${P}_{0}^{H}$, which is the RIPr of ${P}_{1}^{H}$ on $\mathcal{P}_{0}$.
Proposition A.5 provides an example where the above assumption fails to hold. To provide a concrete example where the assumption does hold, consider a one-parameter exponential family of densities written in canonical form
$p_{\theta}(x)=h(x)\exp(\theta T(x)-A(\theta)),$ (52)
where $A:\mathbb{R}\to\mathbb{R}$ is a convex, differentiable function and
$\mathcal{P}_{0}=\{P_{\theta}:\theta\in[a,b]\}\text{ and }\mathcal{P}_{1}=\{P_{\theta}:\theta\in\Theta_{1}\},$ (53)
$\text{for some }-\infty\leqslant a\leqslant b\leqslant\infty,\Theta_{1}\subseteq\Theta\text{ is such that }\Theta_{1}\cap[a,b]=\emptyset.$
Proposition 4.11.
For a one-parameter exponential family of the form (52), the testing problem in (53) satisfies Assumption 4.10.
Note that Assumption 4.10 is not restricted to univariate problems. Consider the multivariate Gaussian testing setup: $\mathcal{P}_{0}=\{N_{d}(\bm{\mu},I_{d}):\mu_{1}\leqslant a,\mu_{2}=\cdots=\mu_{d}=0\}$ and $\mathcal{P}_{1}=\{N_{d}(\bm{\mu},I_{d}):\mu_{1}\geq b\}$, for some $a<b$. The KL divergence between two such distributions is given by $D_{\operatorname{KL}}(N_{d}(\bm{\mu},I_{d}),N_{d}(\bm{\mu}^{\prime},I_{d}))=\frac{1}{2}\|\bm{\mu}^{\prime}-\bm{\mu}\|_{2}^{2}$. It is straightforward to verify that for $\bm{\mu}^{*}=(a,0,\cdots,0)$,
$D_{\operatorname{KL}}(P,N_{d}(\bm{\mu}^{*},I_{d}))=\inf_{P_{0}\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P,P_{0}),\text{ for any }P\in\mathcal{P}_{1}.$
Hence, $\{\hat{p}_{0,n}\}$ is a constant sequence (i.e., for every $n$, the RIPr density $\hat{p}_{0,n}$ is the density of $N_{d}(\bm{\mu}^{*},I_{d})$) and thus Assumption 4.10 holds true.
Theorem 4.12.
Suppose that $X_{1},X_{2},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}H$, for some distribution $H\in\mathcal{P}^{\epsilon}_{1}$ (where $\mathcal{P}^{\epsilon}_{1}$ is as defined in (20) or (21)) and Assumptions 4.2 and 4.10 hold. We consider $\epsilon$ being sufficiently small so that $c^{\prime\prime}_{H}>c^{\prime}_{H}$, where $c^{\prime\prime}_{H},c^{\prime}_{H}$ are solutions of (41) and (42), respectively with $P_{1}$ and $P_{0}$ being replaced by $P_{\theta_{1}(H)}$ and the RIPr of $P_{\theta_{1}(H)}$ on $\mathcal{P}_{0}$, respectively. Then,
$\frac{1}{n}\log R_{n,\epsilon}^{\text{RIPr,plug-in}}\to r^{H,\epsilon}_{\text{RIPr,plug-in}}\text{ almost surely as }n\to\infty,$
where $r^{H,\epsilon}_{\text{RIPr,plug-in}}\geq D_{\operatorname{KL}}(Q_{1,\epsilon}^{H},Q_{0,\epsilon}^{H})-2(\log c^{\prime\prime}_{H}-\log c^{\prime}_{H})\epsilon-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}+(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon\right),$ and $Q_{1,\epsilon}^{H},Q_{0,\epsilon}^{H}$ are distributions with densities $q_{1,\epsilon}^{H},q_{0,\epsilon}^{H}$ as defined in (14) and (15) with $P_{1}$ and $P_{0}$ being replaced by $P_{\theta_{1}(H)}$ and the RIPr of $P_{\theta_{1}(H)}$ on $\mathcal{P}_{0}$, respectively. If we further assume that for $\theta_{1}\in\Theta_{1}$, the bias $b_{\theta_{1}}(\epsilon):=\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}|\theta(H)-\theta_{1}|$ is a real valued function such that $\lim_{\epsilon\to 0}b_{\theta_{1}}(\epsilon)=0$, then we have
$\lim_{\epsilon\to 0}\inf_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}r^{H,\epsilon}_{\text{RIPr,plug-in}}=r^{*},$
where $r^{*}=\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{\theta_{1}},P)$ is the optimal growth rate for testing $\mathcal{P}_{0}$ against $P_{\theta_{1}}$.
The above theorem provides a bound on the growth rate and proves asymptotic optimality as $\epsilon\to 0$, under suitable assumptions. For example, [37, Chapter 4] shows that for estimating the location parameter in a contaminated Gaussian model, the sample median $\hat{\theta}_{n}$ is minimum bias estimator and $b_{\theta_{1}}(\epsilon)=b_{0}(\epsilon)=\Phi^{-1}(\frac{1}{2(1-\epsilon)})$, where $\Phi$ is the Gaussian CDF. Hence $b_{\theta_{1}}(\epsilon)\to\Phi^{-1}(1/2)=0$, as $\epsilon\to 0$, and so it meets all the conditions of the theorem.
5 Simulations
In this section, we present a series of simulations designed to evaluate the performance of our robust tests for both simple and composite hypotheses. We use two key parameters in our analysis: $\epsilon^{A}$, which represents the value of $\epsilon$ specified to the test supermartingale and $\epsilon^{R}$, which denotes the true fraction of data contaminated (A = Algorithm, R = Reality).
5.1 Experiments with simple null
In all the simulation experiments in this subsection, we consider the null $P_{0}$ to be $N(0,1)$, the simple and the composite alternative to be $P_{1}=N(\mu_{1},1)$ for some fixed $\mu_{1}$ and $\mathcal{P}_{1}=\{N(\mu,1):\mu\neq 0\}$ respectively. For simple null vs simple alternative, we employ the test described in Section 2.2. But for $P_{0}$ vs. the composite alternative $\mathcal{P}_{1}$, there is no LFD; hence, we use the test supermartingale introduced in Section 4.1. Both the non-robust predictable plug-in method and our robustified predictable plug-in method for composite alternative, we use the sample median as an estimate of $\mu$. All the results in Fig. 1 and 2 are the average of $10$ independent simulations.
Sanity check under the null.
In this experiment, samples are simulated independently from the following $\epsilon_{R}$-contaminated null distribution: $Q=(1-\epsilon^{R})\times N(0,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$. Here $\epsilon^{A}=\epsilon^{R}=0.01$. This mixture model ensures that the $\epsilon^{R}$ fraction of the sample is drawn from the heavy-tailed Cauchy distribution with location and scale parameters 1-1 and $10$ respectively. For the tests with simple alternative, we consider $\mu_{1}=1$. Fig. 1 illustrates that our robust tests are “safe”, i.e. they do not exhibit growth under the null hypothesis, whereas the non-robust methods show unreliable behaviour with significant fluctuations.
Figure 1: Data is drawn from $(1-\epsilon^{R})\times N(0,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$ and $P_{0}=N(0,1),P_{1}=N(1,1)$, $\epsilon^{A}=\epsilon^{R}=0.01$. Robust tests are safe, but the non-robust tests exhibit unstable and unreliable behavior.
(a) $P_{0}=N(0,1)$, $P_{1}=N(1,1)$
(b) $P_{0}=N(0,1),\mathcal{P}_{1}=\{N(\mu,1):\mu\neq 0\}$
Figure 2: Data is drawn from $(1-\epsilon^{R})\times N(1,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$ and $P_{0}=N(0,1),\mu_{1}=1$, $\epsilon^{A}=\epsilon^{R}=0.1,0.01,0.001$. The growth rate of our robust tests increases as $\epsilon$ decreases. As anticipated, The growth rates for our robust tests based on simple and composite alternatives almost overlap. The growth rates for our robust tests based on simple and composite alternatives in the left and right subfigures look similar.
Growth rate with different contamination.
In this experiment, samples are simulated independently from the mixture distribution $(1-\epsilon^{R})\times N(1,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$ for $\epsilon^{R}=10^{-3},10^{-2},10^{-1}$. This mixture model ensures that the $\epsilon^{R}$ fraction of the sample is drawn from the heavy-tailed Cauchy distribution with location and scale parameters 1-1 and $10$ respectively. For the simple alternative, we consider $\mu_{1}=1$. Fig. 2 shows the growths of processes in logarithmic scale for both simple and composite alternative models: $P_{1}=N(1,1)$ (left) and $\mathcal{P}_{1}=\{N(\mu,1):\mu\neq 0\}$ (right). As expected, The growth rate of our robust tests increases as $\epsilon$ decreases. Notably, both simple and composite tests grow at similar rates (fig. 2 ). It is also evident that non-robust tests exhibit highly erratic behavior, even when plotting the averages of $10$ independent runs.
5.2 Experiments with composite null
In all the simulation experiments in this subsection, we consider the null to be $\mathcal{P}_{0}=\{N(\mu,1):-0.5\leqslant\mu\leqslant 0.5\}$, the simple and the composite alternative to be $P_{1}=N(1,1)$ and $\mathcal{P}_{1}=\{N(\mu,1):\mu\leqslant 0.5\text{ or }\mu\geq 0.5\}$ respectively. For this composite null $\mathcal{P}_{0}$ and simple alternative $P_{1}$, ($N(0.5,1),N(1,1)$) is the LFD pair and we employ the test described in Section 3.2. For this composite null $\mathcal{P}_{0}$ and composite alternative $\mathcal{P}_{1}$, there is no LFD. So, we use our robustified predictable plug-in method for composite alternative, as described in Section 4.3 with the sample median as an estimate of $\mu$. All the results in Fig. 3 and 4 are the average of $10$ independent simulations.
Sanity check under the null.
In this experiment, samples are simulated independently from the following $\epsilon_{R}$-contaminated null distribution: $Q=(1-\epsilon^{R})\times N(0,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$. Here $\epsilon^{A}=\epsilon^{R}=0.01$. For the tests with simple alternative, we consider $\mu_{1}=1$. Fig. 3 illustrates that our robust tests are “safe”, i.e. they do not exhibit growth under the null hypothesis, whereas the non-robust methods show unreliable behavior with significant fluctuations.
Figure 3: Data is drawn from $(1-\epsilon^{R})\times N(0,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$. The null is $\mathcal{P}_{0}=\{N(\mu,1):-0/5\leqslant\mu\leqslant 0/5\}$. Our robust tests are safe, but the non-robust tests exhibit unstable and unreliable behavior.
(a) $P_{1}=N(1,1)$
(b) $\mathcal{P}_{1}=\{N(\mu,1):\mu\leqslant-0.5\text{ or }\mu\geq 0.5\}$
Figure 4: Data is drawn from $(1-\epsilon^{R})\times N(1,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$ and $\epsilon^{A}=\epsilon^{R}=0.1,0.01,0.001$. The growth rate of our robust tests increases as $\epsilon$ decreases. As anticipated, The growth rates for our robust tests based on simple and composite alternatives almost overlap. The growth rates for our robust tests based on simple and composite alternatives in the left and right subfigures look similar.
Growth rate with different contamination.
In this experiment, samples are simulated independently from the mixture distribution $(1-\epsilon^{R})\times N(1,1)+\epsilon^{R}\times\text{Cauchy}(-1,10)$ for $\epsilon^{R}=10^{-3},10^{-2},10^{-1}$. Fig. 4 shows the growths of processes in logarithmic scale for both simple and composite alternative models: $P_{1}=N(1,1)$ (left) and $\mathcal{P}_{1}=\{N(\mu,1):\mu\leqslant-0.5\text{ or }\mu\geq 0.5\}$ (right). As expected, The growth rate of our robust tests increases as $\epsilon$ decreases. Notably, both simple and composite tests grow at similar rates (Fig. 4). It is also evident that non-robust tests exhibit highly erratic behavior, even when plotting the averages of $10$ independent runs.
5.3 Experiment with multivariate data
To evaluate the performance of our proposed robust likelihood ratio test in higher dimensions, we consider $P_{0}=N_{d}(\mathbf{0}_{d},I_{d})$ vs. $P_{1}=N_{d}(\mathbf{1}_{d},I_{d})$, where $\mathbf{0}_{d}$ and $\mathbf{1}_{d}$ are $d$-dimentional vectors whose all elements are 0 and $1$ respectively, $I_{d}$ is the $d\times d$ identity matrix. Data is drawn from $\epsilon^{R}$ contaminated $N_{d}(\mathbf{1}_{d},I_{d})$, where $\epsilon^{R}=0.1$ fraction of the data is corrupted with $d$ dimensional i.i.d. $\text{Cauchy}(-1,10)$ noise. Fig. 5 shows the growths of our processes with $\epsilon^{A}=\epsilon^{R}$ (robust) and SPRT (non-robust) in logarithmic scale for $d=5,10$ and $15$. Our test supermartingale exhibits stable exponential growth, whereas the non-robust SPRT demonstrates highly unstable behavior under contamination. We observe a higher growth rate of our test supermartingale for higher dimensions, which aligns with the fact that the KL divergence (optimal non-robust growth rate) between $N_{d}(\mathbf{0}_{d},I_{d})$ and $N_{d}(\mathbf{1}_{d},I_{d})$ increases with $d$.
Figure 5: Data is drawn from $N_{d}(\mathbf{1}_{d},I_{d})$ contaminated with $d$ dimensional i.i.d. $\text{Cauchy}(-1,10)$. $\epsilon^{R}=\epsilon^{A}=0.1$.
6 Conclusion
In this paper, we established that the likelihood ratio of the least favorable distribution (LFD) pair forms the log-optimal e-value. Extending Huber’s framework, we proved that if an LFD pair exists for a given composite null and alternative, then the LFDs of Huber’s $\epsilon$-contamination neighborhoods around that pair form the optimal LFD pair for the corresponding robustified hypotheses—thereby yielding log-optimal e-values in this broader setting as well. For general composite nulls and alternatives where an exact LFD pair may not exist, we propose new methods and show that our proposed methods are at least asymptotically optimal as $\epsilon\to 0$. Establishing finite $\epsilon$ optimality for tractable testing procedures remains challenging in that setting. We leave this as an avenue for future research.
Acknowledgments
A part of this work was done while AS was a student at the Indian Statistical Institute, Kolkata. We thank Hongjian Wang, Peter Grünwald, Shubhada Agrawal, and Sivaraman Balakrishnan for helpful conversations. We also thank the anonymous reviewers and AE for their helpful suggestions.
Appendix A Proofs
Lemma A.1.
As $\epsilon\downarrow 0$, $c^{\prime\prime}\uparrow\text{ess}\sup_{[\mu]}\frac{p_{1}}{p_{0}}$ and $c^{\prime}\downarrow\text{ess}\inf_{[\mu]}\frac{p_{1}}{p_{0}}$, and therefore,
$q_{1,\epsilon}(X)/q_{0,\epsilon}(X)\to p_{1}(X)/p_{0}(X)\text{ almost surely.}$
Proof of Lemma A.1.
Define,
$f(c)=P_{0}\left[p_{1}/p_{0}<c\right]+\frac{1}{c}P_{1}\left[p_{1}/p_{0}\geq c\right]=1+\int_{p_{1}/p_{0}\geq c}(1/c-p_{0}/p_{1})p_{1}d\mu.$ (54)
Note that $c=c^{\prime\prime}$ is a solution of the equation $f(c)=\frac{1}{1-\epsilon}.$
$f(c+\delta)-f(c)=-\int_{c\leqslant p_{1}/p_{0}\leqslant c+\delta}\left(\frac{1}{c}-\frac{p_{0}}{p_{1}}\right)p_{1}d\mu-\frac{\delta}{c(c+\delta)}\int_{p_{1}/p_{0}\geq c+\delta}p_{1}d\mu$ (55)
Therefore, $\frac{\delta}{c(c+\delta)}\leqslant f(c+\delta)-f(c)\leqslant-\frac{\delta}{c(c+\delta)}P_{1}[p_{1}/p_{0}\geq c+\delta]$ for any $\delta>0$, which implies that $f$ is a continuous and decreasing function.
Let, $c_{0}=\text{ess}\sup_{[\mu]}\frac{p_{1}}{p_{0}}$. If $c_{0}<\infty$, we have $f(c)=1,$ for $c\geqslant c_{0}$ and for $c<c_{0}$, $f(c)$ is strictly decreasing because $f(c+\delta)-f(c)\leqslant-\frac{\delta}{c(c+\delta)}P_{1}[p_{1}/p_{0}\geq c+\delta]<0$, for small $\delta>0$. Therefore, the inverse of $f$ restricted on $0,c_{0}]$ (denoted as $f^{-1}_{[0,c_{0}]}$) is a strictly decreasing and continuous function on the interval $[0,1]$ and hence, $c^{\prime\prime}=f^{-1}_{[0,c_{0}]}(\frac{1}{1-\epsilon})\to f^{-1}_{[0,c_{0}]}(1)=c_{0},$ as $\epsilon\to 0$.
Now, if $c_{0}=\infty$, $f(c+\delta)-f(c)\leqslant-\frac{\delta}{c(c+\delta)}P_{1}[p_{1}/p_{0}\geq c+\delta]<0$, for all $c$ and hence $f$ is strictly decreasing with $\lim_{c\to\infty}f(c)=1$. Therefore, the inverse of $f$ is a strictly decreasing and continuous function on the interval $[0,1)$ and hence, $c^{\prime\prime}=f^{-1}_{[0,c_{0}]}(\frac{1}{1-\epsilon})\to\infty=c_{0},$ as $\epsilon\to 0$.
Thus, we have shown $c^{\prime\prime}\uparrow\text{ess}\sup_{[\mu]}\frac{p_{1}}{p_{0}}$, as $\epsilon\to 0.$ Similarly, one can show that $c^{\prime}\downarrow\text{ess}\inf_{[\mu]}\frac{p_{1}}{p_{0}}$, as $\epsilon\to 0.$
Therefore, $\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}=\frac{p_{1}(X)}{p_{0}(X)}\mathds{1}(c^{\prime}\leqslant\frac{p_{1}(X)}{p_{0}(X)}\leqslant c^{\prime\prime})+c^{\prime}\mathds{1}(\frac{p_{1}(X)}{p_{0}(X)}<c^{\prime})+c^{\prime\prime}\mathds{1}(\frac{p_{1}(X)}{p_{0}(X)}>c^{\prime\prime})\to\frac{p_{1}(X)}{p_{0}(X)},$ $\mu$-almost surely as $\epsilon\to 0.$
Lemma A.2.
Suppose that $D_{\operatorname{KL}}(P_{1},P_{0})<\infty$. Then, $(c^{\prime\prime}-c^{\prime})\epsilon\to 0,$ as $\epsilon\to 0.$
Proof of Lemma A.2.
Let, $c_{0}=\text{ess}\sup_{[\mu]}\frac{p_{1}}{p_{0}}$. If $c_{0}<\infty$, it follows from Lemma A.1 that $c^{\prime\prime}\leqslant c_{0}$ and so $c^{\prime\prime}\epsilon\to 0$, as $\epsilon\to 0$.
Now, if $_{0}=\infty, c′′c^{\prime\prime}\to\infty$ as $\epsilon\to 0.$
From (16), $1+\frac{1}{c^{\prime\prime}}P_{1}\left[p_{1}/p_{0}\geq c^{\prime\prime}\right]\geq\frac{1}{1-\epsilon}$, which implies
$c^{\prime\prime}\epsilon\leqslant(1-\epsilon)P_{1}\left[p_{1}/p_{0}\geq c^{\prime\prime}\right]$ (56)
If $D_{\operatorname{KL}}(P_{1},P_{0})<\infty$, we have $\mathbb{E}_{P_{1}}\left|\log(p_{1}/p_{0})\right|<\infty.$ Then,
$P_{1}\left[p_{1}/p_{0}\geq c^{\prime\prime}\right]=P_{1}\left[\log(p_{1}/p_{0})\geq\log c^{\prime\prime}\right]$ $\leq P_{1}\left[|\log(p_{1}/p_{0})|\geq\log c^{\prime\prime}\right]$
$\leqslant\frac{\mathbb{E}_{P_{1}}|\log(p_{1}/p_{0})|}{\log c^{\prime\prime}}\to 0,$
as $c^{\prime\prime}\to\infty$. Hence, $c^{\prime\prime}\epsilon\to 0,$ since $c^{\prime\prime}\to\infty$, as $\epsilon\to 0$ for the case when $c_{0}=\infty$. ∎
Lemma A.3.
Assume that $P_{i}({p}_{1}/p_{0}=c)=0$, for all $c\in\mathbb{R}$ and for $i=0,1$. If $\hat{p}_{1,n}\to p_{1}$ almost surely as $n\to\infty$ and $c^{\prime}_{n},c^{\prime\prime}_{n}$ are solutions of (33) and (34) respectively, then $c^{\prime\prime}_{n}\to c^{\prime\prime}$ and $c^{\prime}_{n}\to c^{\prime}$ almost surely as $n\to\infty$, where $c^{\prime}$ and $c^{\prime\prime}$ are solutions of (16) and (17).
Proof.
Define, $A_{n}=\{x:\hat{p}_{1,n}(x)/p_{0}(x)>c\}$ and $A=\{x:{p}_{1}(x)/p_{0}(x)>c\}$.
For $x\in A$ (where ${p}_{1}/p_{0}>c$): There exists an $N$ such that for all $n\geq N$, $\hat{p}_{1,n}/p_{0}>c$. Hence, $x\in A_{n}$ for $n\geq N$. This implies that:
$A\subseteq\liminf_{n\to\infty}A_{n}.$
For $x\notin A$ (where ${p}_{1}/p_{0}\leqslant c$):
  • If ${p}_{1}/p_{0}<c$, then for sufficiently large $n$, $\hat{p}_{1,n}<c$, and hence $x\notin A_{n}$.
  • If ${p}_{1}/p_{0}=c$, then the set of such points forms the boundary. By assumption, this set has zero probability.
Thus:
$P_{1}(\limsup_{n\to\infty}A_{n})\leqslant P_{1}(A)\leqslant P_{1}(\liminf_{n\to\infty}A_{n}).$
Consider:
$\hat{p}_{1,n}(A_{n})=\int_{A_{n}}\hat{p}_{1,n}\,d\mu.$
Since $\hat{p}_{1,n}\to{p}_{1}$ pointwise, Scheffe’s theorem gives $\int|\hat{p}_{1,n}-p_{1}|d\mu\to 0.$
$\int_{A_{n}}\hat{p}_{1,n}\,d\mu\leqslant\int_{A_{n}}{p}_{1}\,d\mu+\int_{A_{n}}|\hat{p}_{1,n}-p_{1}|\,d\mu\leqslant\int_{A_{n}}{p}_{1}\,d\mu+\int|\hat{p}_{1,n}-p_{1}|\,d\mu.$
Therefore:
$\limsup_{n\to\infty}\hat{p}_{1,n}(A_{n})\leqslant\limsup_{n\to\infty}\int_{A_{n}}{p}_{1}\,d\mu=\limsup_{n\to\infty}P_{1}(A_{n})\leqslant P_{1}(\limsup_{n\to\infty}A_{n})\leqslant P_{1}(A).$
Similarly.
$\int_{A_{n}}\hat{p}_{1,n}\,d\mu\geq\int_{A_{n}}{p}_{1}\,d\mu-\int_{A_{n}}|\hat{p}_{1,n}-p_{1}|\,d\mu\geq\int_{A_{n}}{p}_{1}\,d\mu-\int|\hat{p}_{1,n}-p_{1}|\,d\mu.$
Therefore:
$\liminf_{n\to\infty}\hat{p}_{1,n}(A_{n})\geq\liminf_{n\to\infty}\int_{A_{n}}{p}_{1}\,d\mu=\liminf_{n\to\infty}P_{1}(A_{n})\geq P_{1}(\liminf_{n\to\infty}A_{n})\geq P_{1}(A).$
Combining the upper and lower bounds, we conclude:
$\lim_{n\to\infty}\hat{p}_{1,n}[\hat{p}_{1,n}/p_{0}>c]={P}_{1}[{p}_{1}/p_{0}>c].$
Similarly, one can show that
$\lim_{n\to\infty}\hat{p}_{1,n}[\hat{p}_{1,n}/p_{0}<c]={P}_{1}[{p}_{1}/p_{0}<c].$
Define,
$f_{n}(c)=P_{0}\left[\hat{p}_{1,n}/p_{0}<c\right]+\frac{1}{c}\hat{P}_{1,n}\left[\hat{p}_{1,n}/p_{0}\geq c\right],\quad f(c)=P_{0}\left[p_{1}/p_{0}<c\right]+\frac{1}{c}P_{1}\left[p_{1}/p_{0}\geq c\right]$
$g_{n}(c)=\hat{P}_{1,n}\left[\hat{p}_{1,n}/p_{0}>c\right]+cP_{0}\left[\hat{p}_{1,n}/p_{0}\leqslant c\right],\quad g(c)={P}_{1}\left[{p}_{1}/p_{0}>c\right]+cP_{0}\left[{p}_{1}/p_{0}\leqslant c\right].$
Then, it follows from what we have shown above that $f_{n}\to f$ and $g_{n}\to g$ pointwise. In Lemma A.1, we have shown that $f_{n},g_{n},f,g$ are all strictly monotone and continuous. Therefore, pointwise convergence implies uniform convergence, and hence, $c^{\prime\prime}_{n}=f_{n}^{-1}(\frac{1}{1-\epsilon})\to f^{-1}(\frac{1}{1-\epsilon})=c^{\prime\prime}$ and $c^{\prime}_{n}=g_{n}^{-1}(\frac{1}{1-\epsilon})\to g^{-1}(\frac{1}{1-\epsilon})=c^{\prime}$. ∎
Proof of Theorem 4.3.
We have $\hat{p}_{1,n}\to p_{1}^{H}$. It follows from Lemma A.3 that $c^{\prime\prime}_{n}\to c^{\prime\prime}_{H},c^{\prime}_{n}\to c^{\prime}_{H}.$ Therefore, for any fixed $\epsilon$, $\hat{q}_{n,1,\epsilon}/\hat{q}_{n,0,\epsilon}\to{q}_{1,\epsilon}^{H}/q_{0,\epsilon}^{H}$ almost surely as $n\to\infty$ and by bounded convergence theorem, $\mathbb{E}_{X\mid X^{n-1}\sim P_{0}}\left[\frac{\hat{q}_{n,1,\epsilon}(X)}{\hat{q}_{n,0,\epsilon}(X)}\mid X^{n-1}\right]\to\mathbb{E}_{X\mid X^{n-1}\sim P_{0}}\left[\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}\mid X^{n-1}\right]$ as $n\to\infty$ and that would immediately imply $\log\hat{E}_{\epsilon,n}(X_{n})-\log E_{\epsilon}^{H}(X_{n})\to 0$ almost surely as $n\to\infty$, where
$E_{\epsilon}^{H}(x)=\frac{\frac{{q}^{H}_{1,\epsilon}(x)}{q^{H}_{0,\epsilon}(x)}}{\mathbb{E}_{X\mid X^{n-1}\sim P_{0}}\left[\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}\mid X^{n-1}\right]+(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon}.$ (57)
Therefore,
$\frac{1}{n}\log R_{n,\epsilon}^{\text{plug-in}}=\frac{1}{n}\sum_{i=1}^{n}\left(\log\hat{E}_{\epsilon,i}(X_{i})-\log E^{H}_{\epsilon}(X_{i})\right)+\frac{1}{n}\sum_{i=1}^{n}\log E_{\epsilon}^{H}(X_{i})\to\mathbb{E}_{H}\log E^{H}_{\epsilon}(X),$
almost surely, since the first term converges to 0 (since $\log\hat{E}_{\epsilon,n}(X_{n})-\log E_{\epsilon}^{H}(X_{n})\to 0$ and hence their average would converge to 0) and the second term converges to $r_{H,\epsilon}^{\text{plug-in}}=\mathbb{E}_{H}\log E^{H}_{\epsilon}(X)$, by SLLN.
Now, one can easily show that
$r_{H,\epsilon}^{\text{plug-in}}=\mathbb{E}_{H}\log E^{H}_{\epsilon}(X)\geq D_{\operatorname{KL}}(Q^{H}_{1,\epsilon},Q^{H}_{0,\epsilon})-2(\log c^{\prime\prime}_{H}-\log c^{\prime}_{H})\epsilon-\log(1+2(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon).$
Proposition A.4.
Assume that (44) holds and $\inf_{P\in\mathcal{P}_{0}}D_{\operatorname{KL}}(P_{1},P)<\infty$. Then, $B_{\epsilon}(x)\to B^{*}(x)$ $\mu$-almost surely as $\epsilon\to 0$. In other words, for any fixed $n\in\mathbb{N}$, $R_{n,\epsilon}^{\text{RIPr}}\to\prod_{i=1}^{n}p_{1}(X_{i})/p_{0}(X_{i})$ $\mu$-almost surely as $\epsilon\to 0.$
Proof.
It is easy to verify that Lemma A.1 and Lemma A.2 holds when $P_{0}$ is sub-probability distribution as well. Now, Lemma A.1 implies $\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\to\frac{p_{1}(X)}{p_{0}(X)}$ almost surely, as $\epsilon\to 0$ and therefore, $\mathbb{E}_{P}\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\to\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}$, as $\epsilon\to 0$. Since $0\leqslant c^{\prime}<c^{\prime\prime}$, Lemma A.2 implies $(c^{\prime\prime}-c^{\prime})\epsilon\to 0,$ since we assumed $D_{\operatorname{KL}}(P_{1},P_{0})<\infty$.
Since $\frac{p_{1}(X)}{p_{0}(X)}$ is an e-variable for $\mathcal{P}_{0}$ [4], we have $\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}\leqslant 1$, for all $P\in\mathcal{P}_{0}$, i.e., $\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}\leqslant 1.$
To show the reverse inequality, define, $B^{\prime}(x):=\frac{\frac{p_{1}(X)}{p_{0}(X)}}{\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}}$. Then it is clear that $\mathbb{E}_{P}B^{\prime}(X)\leqslant 1$ for all $P\in\mathcal{P}_{0}$. Since $\frac{p_{1}(X)}{p_{0}(X)}$ is log-optimal, we have $\mathbb{E}_{P}\log B^{\prime}(X)\leqslant\mathbb{E}_{P}\log\frac{p_{1}(X)}{p_{0}(X)}$, which implies $\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}\geq 1$.
Combining the above two arguments, we obtain $\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{p_{1}(X)}{p_{0}(X)}=1.$
Note that
$\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}=\mathbb{E}_{P_{0}}\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\to\mathbb{E}_{P_{0}}\frac{p_{1}(X)}{p_{0}(X)}=1,$
as $\epsilon\to 0$. Therefore,
$\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]\to 1,\text{ as }\epsilon\to 0$ (58)
Thus, we obtain $B_{\epsilon}(X)\to B^{*}(X)$ almost surely as $\epsilon\to 0$ and for any $n\in\mathbb{N}$,
$R_{n,\epsilon}^{\text{RIPr}}=\prod_{i=1}^{n}B_{\epsilon}(X_{i})\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\prod_{i=1}^{n}B^{*}(X_{i})=\prod_{i=1}^{n}p_{1}(X_{i})/p_{0}(X_{i})\text{ as }\epsilon\to 0.$ (59)
Proof of Proposition 4.5.
We prove it for $\theta_{0}^{*}=\sup\Theta_{0}\leqslant\theta_{1}$ only, and the other case, when $\inf\Theta_{0}\geq\theta_{1}$ can be proved in the same way. The first part that the RIPr is $P_{\theta_{0}^{*}}$ follows from the proof of Proposition 4.11. The exponential family $\{P_{\theta}\}$ with one-dimensional sufficient statistic $T(x)$ satisfies the monotone likelihood ratio (MLR) property in $T(x)$, and since $\frac{p_{\theta_{1}}(x)}{p_{\theta_{0}^{*}}(x)}$ is non-decreasing in $T(x)$, it follows that $\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}=\max\{c^{\prime},\min\{p_{\theta_{0}^{*}}(X)/p_{\theta_{1}}(X),c^{\prime\prime}\}\}$ is also non-decreasing in $T(x)$. It follows from properties of MLR families that $T(X)$ is stochastically increasing in $\theta$. That is, for $X\sim P_{\theta}$ and $X^{\prime}\sim P_{\theta^{*}}$ with $\theta\leqslant\theta^{*}$, we have $P(T(X)>x)\leqslant P(T(X^{\prime})>x)$ for all $x$, which is denoted as $T(X)\preceq_{\text{st}}T(X^{\prime})$ (see e.g., [42, Chapter 3.4]) and hence,
$\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\preceq_{\text{st}}\frac{q_{1,\epsilon}(X^{\prime})}{q_{0,\epsilon}(X^{\prime})}.$
Therefore, it immediately follows that
$\mathbb{E}_{P_{\theta}}\left(\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right)\leqslant\mathbb{E}_{P_{\theta^{*}}}\left(\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right),\forall\theta\in\Theta_{0}.$
Thus, (44) holds. ∎
Proof of Theorem 4.7.
By SLLN,
[…]
where $r^{Q,\epsilon}_{\text{RIPr}}=\mathbb{E}_{Q}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]+(c^{\prime\prime}-c^{\prime})\epsilon\right).$
Note that $D_{\operatorname{TV}}(Q_{1,\epsilon},Q)<2\epsilon$, so
$\left|\mathbb{E}_{Q}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}-\mathbb{E}_{Q_{1,\epsilon}}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right|\leqslant 2(\log c^{\prime\prime}-\log c^{\prime})\epsilon.$ (61)
Therefore,
$r^{\epsilon}_{\text{RIPr}}=\inf_{Q\in H_{1}^{\epsilon}}r^{Q,\epsilon}_{\text{RIPr}}\geq D_{\operatorname{KL}}(Q_{1,\epsilon},Q_{0,\epsilon})-2(\log c^{\prime\prime}-\log c^{\prime})\epsilon-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}+(c^{\prime\prime}-c^{\prime})\epsilon\right).$
Proof of Theorem 4.8.
From (41), we get $(1-\epsilon)(k+\frac{1}{c^{\prime\prime}})\geq k$, which implies $kc^{\prime\prime}\leqslant\frac{1}{\epsilon}-1$. Similarly, from (17), we get $kc^{\prime}\geq\frac{\epsilon}{1-\epsilon}$. Therefore, $(\log c^{\prime\prime}-\log c^{\prime})\epsilon\to 0$, as $\epsilon\to 0$. From Lemma A.1, we have $E_{P_{1}}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\to D_{\operatorname{KL}}(P_{1},P_{0})$, as $\epsilon\to 0$. Note that $D_{\operatorname{TV}}(P_{1},Q)\leqslant\epsilon$, so
$\left|\mathbb{E}_{Q}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}-\mathbb{E}_{P_{1}}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right|\leqslant(\log c^{\prime\prime}-\log c^{\prime})\epsilon.$ (62)
Therefore, $\mathbb{E}_{Q}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\to D_{\operatorname{KL}}(P_{1},P_{0})$, as $\epsilon\to 0$. From Lemma A.2 and (58), we have
$\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]+(c^{\prime\prime}-c^{\prime})\epsilon\right)\to 0.$ Hence,
$r^{Q,\epsilon}_{\text{RIPr}}=\mathbb{E}_{Q}\log\frac{q_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}_{1,\epsilon}(X)}{q_{0,\epsilon}(X)}\right]+(c^{\prime\prime}-c^{\prime})\epsilon\right)\to D_{\operatorname{KL}}(P_{1},P_{0}).$ (63)
Thus, $r^{\epsilon}_{\text{RIPr}}=\inf_{Q\in H^{\epsilon}_{1}}r^{Q,\epsilon}_{\text{RIPr}}\to D_{\operatorname{KL}}(P_{1},P_{0})$, as $\epsilon\to 0.$
Proposition A.5.
Consider the null hypothesis that the mean is at most $\mu$: $\mathcal{P}_{0}=\left\{P\in\mathcal{M}_{1}:\mathbb{E}_{P}[X]\leqslant\mu\right\},$ where $\mathcal{M}_{1}$ denotes the set of all probability measures on $[0,1]$. Let $Q_{n}$ be a sequence of probability measures with densities $q_{n}$ that converge pointwise to a density $q$, corresponding to some measure $Q$. Let $P_{n}^{*}$ denote the reverse information projection (RIPr) of $Q_{n}$ onto $\mathcal{P}_{0}$, and let $p_{n}^{*}$ be its density. Let $P^{*}$ denote the RIPr of $Q$ onto $\mathcal{P}_{0}$, with density $p^{*}$. Then, there exists such a sequence $(Q_{n})$ for which $p_{n}^{*}\not\to p^{*}$ pointwise.
Proof.
It follows from [4][Section 5.2] that the RIPr of $Q_{n}$ on $\mathcal{P}_{0}$ satisfies $\frac{p_{n}^{*}(x)}{q_{n}(x)}=\frac{1}{1+\lambda^{*}_{n}(x-\mu)}$, where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation
$\mathbb{E}_{Q_{n}}\left[\frac{Z-\mu}{1+\lambda^{*}_{n}(Z-\mu)}\right]=0.$ (64)
We can write ${p_{n}^{*}(x)}=\frac{q_{n}(x)}{1+\lambda^{*}_{n}(x-\mu)}$. We are given that $q_{n}(x)\to q(x)$ as $n\to\infty.$ We construct a counterexample to show that the RIPr is not continuous under pointwise convergence of the alternative densities.
Fix $\mu=0.5$, and define $q$ to be the density of a binary distribution:
$q(x)=\begin{cases}\frac{1}{2}&\text{if }x=0,\\ \frac{1}{2}&\text{if }x=1,\\ 0&\text{otherwise}.\end{cases}$
One can easily check that the RIPr of $Q$ onto $\mathcal{P}_{0}$ is itself: $P^{*}=Q$, since $Q$ is an element of $\mathcal{P}_{0}$ (also, $\lambda^{*}=0$ solving (64)).
Now define a sequence of densities $(q_{n})$ as follows:
$q_{n}(x)=\left(1-\frac{1}{n}\right)q(x)+\frac{1}{n}\delta_{n}(x),$
i.e., we add a small mass $\frac{1}{n}$ at $x=n$. Then $q_{n}\to q$ pointwise as $n\to\infty$. Let $Q_{n}$ be the distribution corresponding to $q_{n}$.
Now from (64), we get a quadratic equation $A_{n}\lambda_{n}^{*2}+B_{n}\lambda_{n}^{*}+C=0$, where
$A_{n}$ $=-\mu(1-\mu)(n-\mu),$
$B_{n}$ $=\frac{1}{2}(1-\frac{1}{n})\left[-\mu(n+1-2\mu)+(1-\mu)(n-2\mu)\right]+\frac{1}{n}(n-\mu)(1-2\mu),$
$C_{n}$ $=\frac{1}{2}(1-\frac{1}{n})(1-2\mu)+\frac{1}{n}(n-\mu).$
Therefore, $A_{n}C_{n}\not\to 0$ and hence, $\lambda_{n}^{*}\not\to 0$ as $n\to\infty$. Consequently, $p_{n}^{*}\not\to p^{*}$ pointwise. ∎
Proof of Proposition 4.11.
Let, $P_{\theta^{*}}$ be the RIPr of $P_{\theta_{1}}$ on $\mathcal{P}_{0}=\{P_{\theta}:\theta\in[a,b]\text{ for some }-\infty\leqslant a\leqslant b\leqslant\infty\}$ and $\theta_{1}\notin[a,b]$. Then we shall show that $P_{\theta^{*}}$ has density $p_{\theta^{*}}$, where $\theta^{*}$ is the closest element in $[a,b]$ from $\theta_{1}$.
It is enough to show that $\mathbb{E}_{\theta_{1}}(p_{\theta}/p_{\theta^{*}})\leqslant 1$, for all $\theta\in\theta_{0}$ [4, Theorem 4.7]. Now,
$\mathbb{E}_{\theta_{1}}(p_{\theta}(X)/p_{\theta^{*}}(X))=$ $\mathbb{E}_{\theta_{1}}(\exp\{(\theta_{1}-\theta^{*})T_{k}(X)-A(\theta)+A(\theta^{*})\})$
$=$ $\int\exp\{(\theta_{1}-\theta^{*})T_{k}(x)-A(\theta)+A(\theta^{*})+\theta_{1}T(x)-A(\theta_{1})\}h(x)dx$
$=$ $\exp\{A(\theta_{1}-\theta^{*}+\theta_{1})-A(\theta)+A(\theta^{*})-A(\theta_{1})\}$
Since A is convex, its derivative $A^{\prime}$ is increasing and either $\theta_{1}<\theta^{*}\leqslant\theta$ or $\theta_{1}>\theta^{*}\geq\theta$. So,
$A(\theta_{1}-\theta^{*}+\theta_{1})-A(\theta_{1})$ $=\int_{0}^{1}(\theta-\theta^{*})A^{\prime}(\theta_{1}+t(\theta-\theta^{*}))dt$
$\leqslant\int_{0}^{1}(\theta-\theta^{*})A^{\prime}(\theta_{1}+t(\theta^{*}-\theta^{*}))dt$
$=A(\theta)-A(\theta^{*}).$
Therefore, $\mathbb{E}_{\theta_{1}}(p_{\theta}/p_{\theta^{*}})\leqslant 1$, for all $\theta\in\theta_{0}$.
Since $\hat{\theta}_{n}\to{\theta}_{1}(H)$, $\hat{\theta}^{*}_{n}$ converges to either $a$ or $b$, we call the limit $\theta_{0}(H)$. We have $\hat{\theta}^{*}_{n}\to{\theta}_{0}(H)$. We also have shown above that the RIPr $P_{0}^{H}$ of $P_{1}^{H}$ has density $p_{0}^{H}=p_{{\theta}_{0}(H)}$. So, we now have that $\hat{p}_{0,n}\to p_{0}^{H}$. Therefore, Assumption 4.10 holds. ∎
Proof of Theorem 4.12.
By the above lemma, its RIPr would have density $\hat{p}_{0,n}=p_{\hat{\theta}^{*}_{n}}$, where $\hat{\theta}^{*}_{n}$ is the nearest element in $[a,b]$ from $\hat{\theta}_{n}\notin[a,b]$ (so $\hat{\theta}^{*}_{n}$ is either $a$ or $b$). Since $\hat{\theta}_{n}\to{\theta}_{1}(H)$, $\hat{\theta}^{*}_{n}$ converges to either $a$ or $b$, we call the limit $\theta_{0}(H)$. We have $\hat{\theta}^{*}_{n}\to{\theta}_{0}(H)$. By lemma, we also have that the RIPr $P_{0}^{H}$ of $P_{1}^{H}$ has density $p_{0}^{H}=p_{{\theta}_{0}(H)}$. Therefore, we now have that $\hat{p}_{0,n}\to p_{0}^{H}$, and $\hat{p}_{1,n}\to p_{1}^{H}$. Define,
$B_{\epsilon}^{H}(x)=\frac{\frac{{q}^{H}_{1,\epsilon}(x)}{q^{H}_{0,\epsilon}(x)}}{\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{X\sim P}\left[\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}\right]+(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon}.$
Since $c^{\prime\prime}_{n}\to c^{\prime\prime}_{H},c^{\prime}_{n}\to c^{\prime}_{H}$, we have $\hat{q}_{n,1,\epsilon}/\hat{q}_{n,0,\epsilon}\to{q}^{H}_{1,\epsilon}/q^{H}_{0,\epsilon}$ almost surely as $n\to\infty$ and that would immediately imply $\log\hat{B}_{\epsilon,i}(X_{i})-\log B^{H}_{\epsilon}(X_{i})\to 0$ almost surely as $n\to\infty$.
$\frac{1}{n}\log R_{n,\epsilon}^{\text{plug-in,RIPr}}=\frac{1}{n}\sum_{i=1}^{n}\left(\log\hat{B}_{n,\epsilon}(X_{i})-\log B^{H}_{\epsilon}(X_{i})\right)+\frac{1}{n}\sum_{i=1}^{n}\log B^{H}_{\epsilon}(X_{i})\to r^{H,\epsilon}_{\text{RIPr,plug-in}},$
since the first term converges to 0 and the second term converges to $r^{Q,\epsilon}_{\text{RIPr,plug-in}}=\mathbb{E}_{H}\log B^{H}_{\epsilon}(X)$, by SLLN. Imitating the proof in Theorem 4.7,
$r^{H,\epsilon}_{\text{RIPr,plug-in}}$ $=\mathbb{E}_{H}\log B^{H}_{\epsilon}(X)$
$\geq\mathbb{E}_{P_{\theta_{1}}}\log(q_{1,\epsilon}^{H}/q_{0,\epsilon}^{H})-(\log c^{\prime\prime}_{H}-\log c^{\prime}_{H})\epsilon-\log\left(\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}+(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon\right).$
From (56), $c^{\prime\prime}_{H}\epsilon\leqslant(1-\epsilon)P_{1}^{H}\left[p_{1}^{H}/p_{0}\geq c^{\prime\prime}_{H}\right]\leqslant\frac{\mathbb{E}_{P_{{\theta}_{1}(H)}}|\log(p_{{\theta}_{1}(H)}/p_{\theta_{0}(H)})|}{\log c^{\prime\prime}_{H}}.$
So, $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}c^{\prime\prime}_{H}\epsilon\leqslant\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}\frac{\mathbb{E}_{P_{{\theta}_{1}(H)}}|\log(p_{{\theta}_{1}(H)}/p_{\theta_{0}(H)})|}{\log c^{\prime\prime}_{H}}=0$, since
$\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}\mathbb{E}_{P_{{\theta}_{1}(H)}}|\log(p_{{\theta}_{1}(H)}/p_{\theta_{0}(H)})|<\infty$, as $b_{\theta_{1}}(\epsilon)<\infty$ and
$\lim_{\epsilon\to 0}\inf_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}\log c^{\prime\prime}_{H}=\infty$.
Therefore, $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}(c^{\prime\prime}_{H}-c^{\prime}_{H})\epsilon=0$ and $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}(\log c^{\prime\prime}_{H}-\log c^{\prime}_{H})\epsilon=0.$
Note that $\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}=\min\{c^{\prime}_{H},\max\{c^{\prime\prime}_{H},\exp((\theta_{1}(H)-\theta_{0}(H))T(x)-A(\theta_{1}(H))+A(\theta_{0}(H)))\}\}.$
Since $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}|\theta_{1}(H)-\theta_{1}|=0$, $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}|\theta_{0}(H)-\theta_{0}|=0$, where $\theta_{0}$ is the RIPr corresponding to $\theta_{1}$ (using previous lemma).
Therefore, $\lim_{\epsilon\to 0}\sup_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{{q}^{H}_{1,\epsilon}(X)}{q^{H}_{0,\epsilon}(X)}=\sup_{P\in\mathcal{P}_{0}}\mathbb{E}_{P}\frac{p_{\theta_{1}}(X)}{p_{\theta_{0}}(X)}=1$
Similarly, $\lim_{\epsilon\to 0}\inf_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}\mathbb{E}_{P_{\theta_{1}}}\log(q_{1,\epsilon}^{H}/q_{0,\epsilon}^{H})=\mathbb{E}_{P_{\theta_{1}}}\left(\log\frac{p_{\theta_{1}}(X)}{p_{\theta_{0}}(X)}\right)=D_{\operatorname{KL}}(P_{\theta_{1}},P_{\theta_{0}}).$
Therefore, we proved
$\inf_{H:D_{\operatorname{TV}}(H,P_{\theta_{1}})\leqslant\epsilon}r^{H,\epsilon}_{\text{RIPr,plug-in}}\to r^{*}.$